### v1

# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

Abstract
Elliptically symmetric distributions are a classic example of a semiparametric model where the location vector and the scatter matrix (or a parameterization of them) are the two finite-dimensional parameters of interest, while the density generator represents an infinite-dimensional nuisance term. This basic representation of the elliptic model can be made more accurate, rich, and flexible by considering additional finite-dimensional nuisance parameters. Our aim is therefore to investigate the deep and counter-intuitive links between statistical efficiency in estimating the parameters of interest in the presence of both finite and infinite-dimensional nuisance parameters. Unlike previous works that addressed this problem using Le Cam’s asymptotic theory, our approach here is purely geometric: efficiency will be analyzed using tools such as projections and tangent spaces embedded in the relevant Hilbert space. This allows us to obtain original results also for the case where the location vector and the scatter matrix are parameterized by a finite-dimensional vector that can be partitioned in two sub-vectors: one containing the parameters of interest and the other containing the nuisance parameters. As an example, we illustrate how the obtained results can be applied to the well-known “low-rank” parameterization. Furthermore, while the theoretical analysis will be developed for Real Elliptically Symmetric (RES) distributions, we show how to extend our results to the case of Circular and Non-Circular Complex Elliptically Symmetric (C-CES and NC-CES) distributions.
Index Terms:
Semiparametric models, elliptically symmetric distributions, nuisance parameters, shape matrix, scatter matrix, efficiency, Fisher information matrix, Cramér-Rao bound.
publicationid: pubid: $00$ 2023 IEEE
I Introduction
Asemiparametric model is a statistical model that involves not only a finite-dimensional parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ but also an infinite-dimensional parameter, i.e. a function $g$, that often represents a nuisance parameter. This characterization is general enough to include many well-known examples: the symmetric location model, linear and logistic regression, errors in variables models, missing data and censoring models, copula models and even times series models such as ARMA or ARCH models. We refer the reader to [1, Sect. 4] and [2] for a discussion of how the semiparametric formalism applies to the above-mentioned examples. The clear advantage of adopting a semiparametric model is in the potential gain in term of robustness with rest to some missing knowledge about the random experiment at hand that are indeed required when we use a parametric model. On the other hand, the fact that the function $g$ is left fully unspecified may lead to some efficiency losses in the estimation of the parameter vector of interest $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ in respect of the parametric case. Nevertheless, there are cases in which parametric and semi-parametric efficiency coincide. In other words, in these specific cases, we can gain robustness without losing efficiency. The simplest example is the symmetric location model, in which the goal is to estimate the location parameter without assuming any specific functional form for the (symmetric) density of the data [1, Sect. 3.4, Ex. 1]. In this work, we will analyze the parametric and the semiparametric efficiency for the estimation of the parameter vector of interest when the considered semiparametric model contains nuisance parameters of both finite and infinite dimension. Specifically, we will focus on the statistical model of the elliptically symmetric distributions.
Elliptically symmetric distributions have established themselves as a statistical model capable of capturing the heterogeneous nature of data in a wide range of applications: from remote sensing and communication to seismology and finance (see e.g. [3, 4, 5, 6, 7] for a complete list of references and examples). As for a Gaussian distribution, an elliptical distribution has the advantage of being fully characterized by the location vector $\bm{\mu}$ and the covariance/scatter matrix $\bm{\Sigma}$, while its flexibility with respect to (w.r.t.) the latter is provided by the density generator, i.e. a function $g$, that is able to characterizes the heavy or light-tailed behavior of the data. To this end, the density generator $g$ may depend on additional parameters that control the “level of the tails” of the resulting distribution. Among the most popular and widely-used elliptical distributions, we may cite the Student $t$-distribution (characterized by the, so called, degree of freedom $\lambda$) and the Generalized Gaussian distribution (characterized by the scale and shape parameters) [8].
Inference in elliptical distributions generally requires the estimation of $\bm{\mu}$ and $\bm{\Sigma}$ as main parameters of interest in the eventual presence of additional nuisance terms. For clarity, we would like to recall here that a nuisance term is an unknown parameter whose estimation is not strictly required but that can have an impact on the estimation performance, i.e. the efficiency, of the parameters of interest.
To formalize this inference problem, we need to specify the statistical model we are considering. Three options, ranging from the most to the least restrictive, can be considered:
  • M1
    Parametric modeling with fully specified density generator $g$. This is the less general case in which a full knowledge (i.e. both its functional form and its parameters) of the specific density generator $g$ is assumed to be a-priori known. To fix the idea, in order to derive estimation procedure for $\bm{\mu}$ and $\bm{\Sigma}$, a practitioner may assume to know a-priori that the data follow a $t$-distribution with an a-priori known degree of freedom $\lambda$. This full knowledge of the density generator is not realistic in practice and need to be relaxed.
  • M2
    Parametric modeling with specified density generator $g$ up to its parameters. This is the classical approach in parametric elliptical inference and most of the literature deals with this case. Returning once again to the example of the $t$-distribution, in this case a practitioner may suppose a-priori that the data generating process is a $t$-distribution characterized by the unknown degree of freedom $\lambda$. Consequently, unlike in M1, the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$ need to be estimated together with the nuisance term $\lambda$. In general this lack of knowledge on the true value of $\lambda$ may lead to a performance degradation, i.e. efficiency losses, in the estimation of $\bm{\mu}$ and $\bm{\Sigma}$. Even if more flexible w.r.t. the previous modeling approach, the requirement of the a-priori knowledge of the functional form of the density generator $g$ may be questionable. This leads to the semiparametric modeling.
  • M3
    Semiparametric modeling where the functional form of the density generator $g$ is left unspecified. This is the most realistic case in which the practitioner only supposes that the data are elliptically distributed, without assuming any specific density generator $g$, while estimating the parameters of interest $\bm{\mu}$ and $\bm{\Sigma}$. As a consequence, the density generator has to be considered as an infinite-dimensional nuisance term. Intuitively, one might expect efficiency losses in this case to be greater than those incurred in M2. After all, in model M3, it is the entire functional form of $g$ that is unknown, not just the value of theirs characterizing parameters.
As this discussion suggests, it would be useful to carefully analyze the efficiency losses in the estimation of the parameters of interest when moving from the more to the less stringent modeling. In particular, the following question may arise: is it possible to relax unrealistic assumptions on the adopted statistical model (for example, moving from M2 to M3, or even from M1 to M3) without losing efficiency?
As showed by Hallin and Paindaveine in two seminal papers [9, 10], the answer may be surprising and counter-intuitive. They proved that a decisive role is played by the additional finite-dimensional nuisance parameters that are involved, in an implicit or explicit manner, in the semiparametric modeling of elliptical distributions. Specifically, in their works, Hallin and Paidaveine used the Le Cam’s asymptotic theory to bring to light the fundamental role of a finite-dimensional nuisance parameter hidden in the semiparametric elliptic model: the scale parameter. In short, it is well-known that, in an elliptical model, the scatter matrix $\bm{\Sigma}$ and the density generator $g$ are not jointly identifiable due to a scale ambiguity. In order to remove this ambiguity, $\bm{\Sigma}$ must be rewritten as the product of a scale parameter $s\triangleq S(\bm{\Sigma})$ and of a shape matrix $\mathbf{V}_{S}\triangleq\bm{\Sigma}/S(\bm{\Sigma})$, where $S(\cdot)$ is a given homogeneous functional. In the resulting inference problem, $\bm{\mu}$ and $\mathbf{V}_{S}$ are to be considered as parameters of interest, while $s$ and $g$ are finite and infinite-dimensional nuisance parameters, respectively. Then, in [9, 10], it has been shown that, if the scale parameter $s$ is considered as a nuisance term, then: i) not knowing the function form of $g$ does not lead to any efficiency loss on the estimation of $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which $g$ is supposed to be a-priori known, and ii) if a determinant-based scale function $S(\cdot)$ is adopted, then knowing $g$ and $s$ does not lead to efficiency losses on $\bm{\mu}$ and $\mathbf{V}_{S}$ w.r.t. the case in which both $g$ and $s$ are perfectly known.
Inspired by this surprising result, in the first part of this paper, we focus on the three models M1, M2 and M3 discussed above (that are slightly different from the scenario analyzed in [9, 10]). Unlike [9, 10] where the main analytical tool is Le Cam’s asymptotic theory, we will use a different approach to the analysis of semiparametric models. Specifically, we will follow the geometric, Hilbert-space-based approach developed in the foundational monograph [1], as it has the potential to reach a broader audience. The main rationale underlying this geometrical approach is that the structure of an Hilbert space enables the unified treatment of both finite-dimensional and infinite-dimensional parameters.
To make this work as self-contained as possible, in Sect. II we present the geometrical tools (i.e. projections and tagent spaces) that will be at the core of our analysis of the semiparametric elliptical model. Moreover, three fundamental Lemmas are provided and their proofs detailed in the Appendix. We move then to Sect. III where, after a brief recall of the main definitions and properties of the Real Elliptically Symmetric (RES) distributions, we provide an extensive discussion of score vectors and related Fisher Information Matrix (FIM) for the parametric RES model. In doing this, we will make use of some fundamental outcomes obtained in [9, 10] for the inference of the location $\bm{\mu}$ and the shape $\mathbf{V}_{S}$ in the presence of the nuisance scale parameter $s$. Furthermore, we introduce the original analysis of an equivalent RES model, important in many applications, parameterized by $\bm{\mu}$ and the constrained covariance/shape matrix $\bm{\Sigma}$. In Sect. IV the geometrical tools introduced in Sect. II will be specified for the semiparametric RES elliptical models under considerations, while our main results of the first part of the paper is collected in Sect. V and in particular in Proposition V.2, where the closed-form expressions of parametric and semiparametric FIM and the related information bounds for the models in M1, M2 and M3 are provided. Proposition V.2 plays a crucial role in formalizing and fully understanding the relationship between the (lack of) efficiency losses and the finite and infinite-dimensional nuisance terms involved in the considered elliptical parametric and semiparametric models.
The second and last part of our paper, basically concentrated in Sect. VI, deals with parameterized elliptical distributions. In particular, we suppose that the location vector $\bm{\mu}(\bm{\theta})$ and the covariance/scatter matrix $\bm{\Sigma}(\bm{\theta})$ are parameterized by a finite dimensional vector $\bm{\theta}=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}$, where the sub-vector $\bm{\gamma}$ contains the parameters of interest, while $\bm{\xi}$ collects the finite-dimensional nuisance terms. We aim then to investigate the efficiency losses on the following two scenarios:
  • E1
    Parametric modeling: estimation of $\bm{\gamma}$ in the presence of the finite-dimensional nuisance $\bm{\xi}$ with fully specified density generator $g$. Again, this is the classical parametric context adopted in the vast majority of the applications. Nevertheless, as said before, the assumption of a perfect a-priori knowledge of the density generator $g$ may be unrealistic in practice.
  • E2
    Semiparametric modeling: estimation of $\bm{\gamma}$ in the presence of the finite-dimensional nuisance $\bm{\xi}$ and the infinite-dimensional nuisance $g$. This is the more realistic case in which we only need to assume that the data generating process follows an elliptical distribution while its density generator is considered as an infinite-dimensional unknown term.
So we ask ourselves the same question as before: is it possible to relax the parametric assumption, moving from E1 to E2, without losing efficiency? The answer to this question depends on the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ at hand. Proposition VI.1, which represents the main result of Sect. VI, provides a condition to determine whether the given parameterization leads to efficiency losses or not. As an example, this condition will be applied to two cases, important in many applications, that are i) the parameterization $=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\mapsto(\bm{\mu}(\bm{\gamma}),\bm{\Sigma}(\bm{\xi}))$, i.e. when the location and the scatter matrix have no parameters in common and the ii) “low-rank” parameterization. Then, Sect. VII generalizes all the results previously obtained in the context of RES distributions to the case of Circular Complex Elliptically Symmetric (C-CES) and Non Circular CES (NC-CES) distributions. Finally Sect. VIII concludes the paper and the technical proofs are reported in the Appendix.
Notation: Italics indicates scalar quantities ($a$), lower case and upper case boldface indicate column vectors ($\mathbf{a}$) and matrices ($\mathbf{A}$). The superscripts $T$, *, $H$ and ##{\#}# indicate the transpose, the complex conjugate, the Hermitian and the Moore-Penrose inverse operators such that $\mathbf{A}^{-T}\triangleq(\mathbf{A}^{-1})^{T}=(\mathbf{A}^{T})^{-1}$, $\mathbf{A}^{-*}\triangleq(\mathbf{A}^{-1})^{*}=(\mathbf{A}^{*})^{-1}$, $\mathbf{A}^{-H}\triangleq(\mathbf{A}^{-1})^{H}=(\mathbf{A}^{H})^{-1}$ and $\mathbf{A}^{\#T}\triangleq(\mathbf{A}^{\#})^{T}=(\mathbf{A}^{T})^{\#}$. Moreover, $\mathbf{A}^{-1/2}\triangleq(\mathbf{A}^{-1})^{1/2}$ denotes any square root of the inverse of the symmetric positive definite matrix $\mathbf{A}$. Each entry of a vector $\mathbf{a}$ and of a matrix $\mathbf{A}$ is indicated as $a_{i}\triangleq[\mathbf{a}]_{i}$ and $a_{ij}\triangleq[\mathbf{A}]_{ij}$, respectively. The symbol $\mathrm{vec}$ indicates the standard vectorization operator that maps column-wise the entry of an $m\times m$ matrix $\mathbf{A}$ in an $m^{2}$-dimensional column vector $\mathrm{vec}(\mathbf{A})$. The Hadamard product $\mathbf{B}\odot\mathbf{C}$ is the matrix whose $(i,j)$-th element is $[\mathbf{B}]_{ij}[\mathbf{C}]_{ij}$. The Kronecker product ${\mathbf{B}}\otimes{\mathbf{C}}$ denotes the block matrix whose $(i,j)$ block element is $b_{ij}{\mathbf{C}}$, and the commutation matrix ${\mathbf{K}}_{m}$ is such that $\mathrm{vec}(\mathbf{A}^{T})={\bf K}_{m}\mathrm{vec}(\mathbf{A})$. The operator $\underline{\mathrm{vec}}(\mathbf{A})$ defines the $m^{2}-1$-dimensional vector obtained from $\mathrm{vec}(\mathbf{A})$ by deleting its first element, i.e. $\mathrm{vec}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vec}}(\mathbf{A})^{T}]^{T}$. For any $m\times m$ symmetric matrix $\mathbf{A}$, $\mathrm{vecs}(\mathbf{A})$ indicates the $m(m+1)/2$-dimensional vector of the entries of the lower triangular part of $\mathbf{A}$. The duplication matrix $\mathbf{D}_{m}$ is implicitly defined as the unique $m^{2}\times m(m+1)/2$ matrix satisfying $\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A})$ for any symmetric matrix $\mathbf{A}$ [11, 12]. Let us now implicitly define the operator $\underline{\mathrm{vecs}}(\cdot)$ as $\mathrm{vecs}(\mathbf{A})\triangleq[a_{11},\underline{\mathrm{vecs}}(\mathbf{A})^{T}]^{T}$. The identity matrix of dimension $m$ is indicated as $\mathbf{I}_{m}$ and $\mathbf{e}_{k}$ indicates the $k$th vector of the canonical basis of $\mathbb{R}^{m^{2}}$. Moreover, let $\underline{\mathbf{I}}_{m}$ be the operator such that $\underline{\mathrm{vecs}}(\mathbf{A})=\underline{\mathbf{I}}_{m}\mathrm{vecs}(\mathbf{A})$ that can be obtained from $\mathbf{I}_{m(m+1)/2}$ by removing its first row.
Let $\mathbf{A}(\bm{\theta})$ be a matrix (or possibly a vector or even a scalar) function of the real vector $\bm{\theta}\in\Theta\subseteq\mathbb{R}^{d}$, then $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\theta}_{0})$ while $\mathbf{A}_{i}^{0}\triangleq{\frac{\partial\mathbf{A}(\bm{\theta})}{\partial\theta_{i}}}|_{\bm{\theta}=\bm{\theta}_{0}}$, where $\bm{\theta}_{0}$ is a particular (or true) value of $\bm{\theta}\in\Theta$. Similarly, the gradient of a function $f(\bm{\theta})$ evaluated at $\bm{\theta}_{0}$, i.e. $\nabla_{\bm{\theta}}f(\bm{\theta})|_{\bm{\theta}=\bm{\theta}_{0}}$, will be indicated as $\nabla_{\bm{\theta}}f(\bm{\theta}_{0})$.
[…]
Let us now introduce the Hilbert space $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$ as the (infinite-dimensional) linear space of the $\mathfrak{B}(\mathcal{X})$-measurable scalar functions with zero-mean and finite variance:
$\mathcal{H}\triangleq\left\{f:\mathcal{X}\rightarrow\mathbb{R}|{E_{0}\{h\}}=0,{E_{0}\{h^{2}\}}<+\infty\right\},$ (1)
endowed with the canonical inner product
[…]
We note that the norm associated to the inner product in (2) is $||h||_{\mathcal{H}}=\sqrt{{E_{0}\{h^{2}\}}}$ that is the standard deviation of $h\in\mathcal{H}$.
Let us now introduce the $q$-replicating Hilbert space $\mathcal{H}_{q}=\mathcal{H}\times\cdots\times\mathcal{H}$ as the linear space of the $\mathfrak{F}$-measurable, $q$-variate functions $\mathbf{h}:\mathcal{X}\rightarrow\mathbb{R}^{q}$. This set can clearly be obtained as $q$ Cartesian products of $\mathcal{H}$ in (1) and this explain the name of $q$-replicating space.
Remark: It is worth noticing that the $q$-replicating space $\mathcal{H}_{q}$ has been introduced only to simplify some notation and to improve the readability of some result proposed in the following sections. However, it does not introduce any additional geometrical structure since it is just composed of $q$ copies of $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$. In particular, when we work on (subspaces of) $\mathcal{H}$, instead of $\mathcal{H}_{q}$, all the operators, as expectations and projections, has to be interpreted component-wise (see e.g. [13, Sect. 2.4, Remark 2] or [1, Sect. 2.4]) as it is proved in Lemma .1 and Theorem .1 reported in Appendix .1. Specifically, for each $\mathbf{h}=(h_{1},\ldots,h_{q})^{T}\in\mathcal{H}_{q}$ s.t. $h_{i}\in\mathcal{H}$ and $u\in\mathcal{U}\subseteq\mathcal{H}$, we have:
$[\Pi(\mathbf{h}|\mathcal{U})]_{i}\triangleq\Pi([\mathbf{h}]_{i}|\mathcal{U})=\Pi(h_{i}|\mathcal{U}),$ (3)
$[{E_{0}\{\mathbf{h}u\}}]_{i}\triangleq{E_{0}\{[\mathbf{h}]_{i}u\}}=\left\langle h_{i},u\right\rangle_{\mathcal{H}},$ (4)
for $i=1,\ldots,q$.
II Models, tangent spaces and score vectors
A semiparametric model is a set of probability density function (pdf) parameterized by a finite dimensional vector of interest and by an infinite-dimensional parameter, i.e. a function, that generally plays the role of a nuisance term. The study of the estimation efficiency in semiparametric models, along with the derivation of the relevant information bounds, is a well-established topic in the statistical literature (see e.g. [1, 13, 14] and the reference therein). We refer the reader to [15, 16] for a tutorial introduction on this subject. In this work, we focus on the case in which the considered semiparametric model involves an additional finite dimensional nuisance vector, along with the infinite-dimensional nuisance term. More formally, let us consider the general semiparametric model:
$\mathcal{P}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,g\in\mathcal{G}\right\},$ (5)
[…]
By using a self-explanatory notation, we now introduce three parametric submodels of $\mathcal{P}$ as:
$\mathcal{P}_{1}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi}_{0},g_{0}):\bm{\gamma}\in\Gamma\right\},$ (6)
$\mathcal{P}_{2}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi},g_{0}):\bm{\xi}\in\Psi\right\},$ (7)
$\mathcal{P}_{1,2}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g_{0}):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\},$ (8)
along with the non-parametric model
$\mathcal{P}_{3}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi}_{0},g):g\in\mathcal{G}\right\}.$ (9)
Let us define the score vector $\mathbf{s}_{\bm{\gamma}_{0}}$ in $\mathcal{P}_{1}$ as:
$\left[\mathbf{s}_{\bm{\gamma}_{0}}\right]_{i}\triangleq\left[\mathbf{s}_{\bm{\gamma}_{0}}(\mathbf{x})\right]_{i}=\left.\partial\ln p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi}_{0},g_{0})/\partial\gamma_{i}\right|_{\bm{\gamma}=\bm{\gamma}_{0}},$ (10)
for $i=1,\ldots,q$, that represents the score vector of the parameters of interest. Similarly, the score vector $\mathbf{s}_{\bm{\xi}_{0}}$ of the finite-dimensional nuisance parameters $\bm{\xi}_{0}$ in $\mathcal{P}_{2}$ is given by:
$\left[\mathbf{s}_{\bm{\xi}_{0}}\right]_{j}\triangleq\left[\mathbf{s}_{\bm{\xi}_{0}}(\mathbf{x})\right]_{j}=\left.\partial\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},\bm{\xi},g_{0})/\partial\xi_{j}\right|_{\bm{\xi}=\bm{\xi}_{0}},$ (11)
for $j=1,\ldots,r$. Under the regularity conditions discussed in [17, Sects. 6.2, 6.3], it is immediate to verify that $[\mathbf{s}_{\bm{\gamma}_{0}}]_{i},[\mathbf{s}_{\bm{\xi}_{0}}]_{j}\in\mathcal{H}$, $\forall i,j$, i.e. they have zero-mean and finite variance.
We can now introduce the (finite-dimensional) tangent space of the parametric submodels $\mathcal{P}_{2}$ as the linear span of $\mathbf{s}_{\bm{\xi}_{0}}$ in $\mathcal{H}$ ([13, Sect. 2.3], [16, eq. (5.24)]):
$\mathcal{H}\supseteq\mathcal{T}_{2}\triangleq\mathrm{Span}\{[\mathbf{s}_{\bm{\xi}_{0}}]_{1},\ldots,[\mathbf{s}_{\bm{\xi}_{0}}]_{r}\}.$ (12)
Finally, the nuisance tangent space $\mathcal{T}_{3}\subseteq\mathcal{H}$ of the non-parametric model $\mathcal{P}_{3}$ is defined as in [1, Sect. 3.2, Def. 2], [13, Sect. 4.4]. Note that, by construction, $\mathcal{T}_{2}$ and $\mathcal{T}_{3}$ are finite- and infinite-dimensional closed subspaces of $\mathcal{H}$.
According to the previous definition, let us define the efficient score vector $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ for the vector $\bm{\gamma}_{0}$ of the parameter of interest in the parametric submodel $\mathcal{P}_{1,2}$ in (8) as ([13, Sect. 3.4], [1, Sect. 2.4] and [16, Def. 4]):
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\triangleq\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}),$ (13)
where the projection $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})$ is to be interpreted component-wise as indicated in (3). Since $\mathcal{T}_{2}$ is a finite-dimensional subspace of $\mathcal{H}$, the projection operator $\Pi(\cdot|\mathcal{T}_{2})$ can be derived in closed form as [16, eq. (7)] [13, Sect. 2.4, Ex. 1]:
$\Pi(h|\mathcal{T}_{2})=E_{0}\{h\mathbf{s}_{\bm{\xi}_{0}}^{T}\}\mathbf{I}_{\bm{\xi}_{0}}^{-1}\mathbf{s}_{\bm{\xi}_{0}},\;h\in\mathcal{H},$ (14)
where:
$\mathbf{I}_{\bm{\xi}_{0}}=E_{0}\{\mathbf{s}_{\bm{\xi}_{0}}\mathbf{s}_{\bm{\xi}_{0}}^{T}\},$ (15)
is the FIM for $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{2}$ in (7). Then, we have that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ can be explicitly expressed as:
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-E_{0}\{\mathbf{s}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\xi}_{0}}^{T}\}\mathbf{I}_{\bm{\xi}_{0}}^{-1}\mathbf{s}_{\bm{\xi}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}\mathbf{s}_{\bm{\xi}_{0}},$ (16)
where
$\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\xi}_{0}}^{T}\}$ (17)
is the matrix of the cross-information terms in the parametric submodel $\mathcal{P}_{1,2}$ in (8).
Using the same geometrical approach, the semiparametric efficient score vector $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$ for $\bm{\gamma}_{0}$ in the semiparametric model $\mathcal{P}$ in (5) is given by [1, Sect. 3.4, eq. (18)]:
$\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\triangleq\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}+\mathcal{T}_{3}),$ (18)
where:
$\mathcal{H}\supseteq\mathcal{T}_{2}+\mathcal{T}_{3}\triangleq\left\{h\in\mathcal{H}|h=o+l,o\in\mathcal{T}_{2},l\in\mathcal{T}_{3}\right\},$ (19)
and since $\mathcal{T}_{2}$ is a (closed) finite-dimensional subspace and $\mathcal{T}_{3}$ is closed, then $\mathcal{T}_{2}+\mathcal{T}_{3}$ is closed. In general, since $\mathcal{T}_{2}+\mathcal{T}_{3}$ is infinite-dimensional, a closed form for the projection operator $\Pi(\cdot|\mathcal{T}_{2}+\mathcal{T}_{3})$ does not exist. Fortunately, some further manipulation is still possible. In fact, let us first recall that, for two orthogonal closed subspaces $\mathcal{A}$ and $\mathcal{B}$ of $\mathcal{H}$ we have the following property:
$\Pi(h|\mathcal{A}+\mathcal{B})=\Pi(h|\mathcal{A})+\Pi(h|\mathcal{B}),\;\forall h\in\mathcal{H}.$ (20)
Moreover, it can be noted that $\mathcal{T}_{2}+\mathcal{T}_{3}$ can be expressed as the (direct) sum of the two orthogonal subpsaces $\mathcal{T}_{2}$ and $((\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\bot})$. Then, from (20), we immediately have that:
$\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$$\triangleq$ $\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}+\mathcal{T}_{3})$ (21)
$=$ $\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp})$
$=$ $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}-\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}),$
where the last equality comes from (13) and from the fact that $\mathcal{T}_{2}\supset\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})\perp(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}$.
The relation (21) between the efficient score vectors $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$ and $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ is the key tool to compare the efficient Semiparametric FIM (SFIM)
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})\triangleq E_{0}\left\{\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\bar{\mathbf{s}}_{\bm{\gamma}_{0}}^{T}\right\}$ (22)
for the parameter of interest $\bm{\gamma}_{0}$ in the presence of both the finite- and infinite dimensional nuisance parameters $\bm{\xi}_{0}$ and $g_{0}$ with the efficient FIM
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})\triangleq E_{0}\left\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}^{T}\right\}$ (23)
for $\bm{\gamma}_{0}$ in the presence of only the finite-dimensional nuisance $\bm{\xi}_{0}$, while $g_{0}$ is known. Specifically, for this comparison, we will make use of the following lemma, given without proof in [1, Sect. 3.4, Prop. 3] and of which we provide a full proof in the Appendix 153 of this work.
Lemma II.1.
The efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ and the efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ for $\bm{\gamma}_{0}$ in the presence of respectively, the finite- and infinite-dimensional nuisance terms $\bm{\xi}_{0}$ and $g_{0}$, and only the finite-dimensional nuisance term $\bm{\xi}_{0}$, are connected through the relation:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})-E_{0}\left\{\mathbf{p}\mathbf{p}^{T}\right\},$ (24)
where
$\mathbf{p}\triangleq\Pi\left(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}\right).$ (25)
It is worth noticing that the matrix $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ is the efficient FIM for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance vector $\bm{\xi}_{0}$ in the parametric submodel $\mathcal{P}_{1,2}$ in (8). Consequently the CRB for $\bm{\gamma}_{0}$ in the presence of $\bm{\xi}_{0}$ is given by:
$\mathrm{CRB}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})^{-1}=\left[\mathbf{I}_{\bm{\gamma}_{0}}-\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}^{T}\right]^{-1},$ (26)
where:
$\mathbf{I}_{\bm{\gamma}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\gamma}_{0}}^{T}\},$ (27)
is the FIM for $\bm{\gamma}_{0}$ in the parametric submodel $\mathcal{P}_{1}$ in (6). It is immediate to recognize in (26) the well-known expression of the CRB for parametric estimation in the presence of a finite-dimensional nuisance vector.
Moreover, we note that this lemma implies that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})\leq\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$. This means that the information about the parameter $\bm{\gamma}_{0}$ is reduced or remains the same in the presence of the infinite-dimensional nuisance function $g_{0}$. It is clear from Lemma II.1 that there is no loss of information iff $\mathbf{p}=\mathbf{0}$. The following lemma (also given without proof, in [1, Sect. 3.4, Prop. 3]) and proven in the Appendix .3 of this work, will specify the condition for which the presence of the nuisance function $g_{0}$ will not bring any loss of information.
Lemma II.2.
The efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ for the model $\mathcal{P}$ in (5) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in $\mathcal{P}_{1,2}$, i.e. $\mathbf{p}=\mathbf{0}$ if and only if (iff) the following condition is satisfied
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\triangleq\ \mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})\perp\mathcal{T}_{3}.$ (28)
In the following sections, we will make extensive use of Lemmas II.1 and II.2 to bring to light the sometimes surprising and counter-intuitive relationships between parametric and semiparametric efficiency in the family of elliptical distributions.
III A short introduction to the RES distributions
In this section, we briefly recall the definition and the main properties of the RES distributions distributions. We will then present and discuss a “canonical parameterization” that can be built upon them. Many different, yet equivalent, representation of the RES family can be found in the literature [18, 19, 9, 10, 8]. Here, we adopt the approach and the notation introduced in [8]. Moreover, in the following, we will consider only the absolutely continuous case, i.e. we suppose that each distribution admits a density w.r.t. the Lebesgue measure on $\mathbb{R}^{m}$. Before moving on, we would like to underline that some of the outcomes discussed in the following sections have already been presented in [9, 10] by using a “semiparametric generalization” of the Le Cam theory on Local Asymptotically Normal (LAN) families of distributions [20], [21, Ch.6]. Our main goal here is to recast the problem of the statistical inference in RES distributions in the framework of the Hilbert spaces and present a systematic analysis based on purely geometrical concepts, such as those used in Lemmas II.1 and II.2.
III-A Essentials on RES distributions
A real-valued, random observation vector $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{m}$ is said to be elliptically symmetric distributed if its probability density function (pdf) can be expressed (in the absolutely continuous case) as:
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},g)=|\bm{\Sigma}|^{-1/2}g\left((\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right),$ (29)
where $\bm{\mu}\in\mathbb{R}^{m}$ is a location vector, $\bm{\Sigma}\in\mathcal{S}_{m}^{\mathbb{R}}$ is an $m\times m$, positive definite, scatter matrix in the set $\mathcal{S}_{m}^{\mathbb{R}}$ of the symmetric real matrices.  [^1] In this article we will limit ourselves to considering scatter, covariance and shape matrices as elements of the linear subspace of symmetric matrices and not as elements of the manifold of positive definite matrices. The density generator $g\in\mathcal{G}$ is a function belonging to a set $\mathcal{G}$ such that:
[…]
where the value of $\delta_{m}$ is such that (29) is a proper density that integrates to 1. In the following, the notation $\mathbf{x}\sim RES_{m}(\bm{\mu},\bm{\Sigma},g)$ indicates that a random vector $\mathbf{x}\in\mathcal{X}$ has the density given in (29).
A fundamental result for RES distributed vectors is the Stochastic Representation Theorem. Specifically, if $\mathbf{x}\sim RES_{m}(\bm{\mu},\bm{\Sigma},g)$ then it can be expressed as:
$\mathbf{x}=_{d}\bm{\mu}+\sqrt{\mathcal{Q}}\bm{\Sigma}^{1/2}\mathbf{u},$ (31)
where the random vector $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ is uniformly distributed on the unit sphere $S_{\mathbb{R}}^{m-1}\triangleq\{\mathbf{u}\in\mathbb{R}^{m}|||\mathbf{u}||=1\}$ and consequently satisfies $E\{\mathbf{u}\}=\mathbf{0}$ and $E\{\mathbf{u}\mathbf{u}^{T}\}=m^{-1}\mathbf{I}_{m}$. The positive random variable $\mathcal{Q}$, called 2nd-order modular variate, is such that (s.t.)
$\mathcal{Q}=_{d}Q_{\bm{\mu},\bm{\Sigma}}(\mathbf{x})\triangleq(\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu}),\;\mathbf{x}\in\mathcal{X}$ (32)
and it is independent of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$. Moreover, $\mathcal{Q}$ has pdf given by:
$p_{\mathcal{Q}}(q)=\delta_{m}^{-1}q^{m/2-1}g(q).$ (33)
It is immediate to verify that the definition of elliptical density suffers from a lack of indentifiability for the couple $(\bm{\Sigma},g)$. Specifically, we can easily note that $RES_{m}(\bm{\mu},\bm{\Sigma},g(t))\equiv RES_{m}(\bm{\mu},c^{m/2}\bm{\Sigma},g(ct)),\forall c>0$. To avoid this ambiguity, we decided to put a constraint on the “functional form” of the density generator $g$. In particular, we force $g$ to belong to the following set:
[…]
Note that, from the stochastic representation (31), the properties of $\mathbf{u}\sim S_{\mathbb{R}}^{m-1}$ and the fact that $\mathcal{Q}$ is independent of $\mathbf{u}$, the constraint $E\{\mathcal{Q}\}=m$ implies that
$E\{(\mathbf{x}-\bm{\mu})(\mathbf{x}-\bm{\mu})^{T}\}=E\{\mathcal{Q}\}\bm{\Sigma}^{1/2}E\{\mathbf{u}\mathbf{u}^{T}\}\bm{\Sigma}^{T/2}=\bm{\Sigma}.$ (35)
i.e., the scatter matrix can be directly interpreted as the usual covariance matrix.
Let us now introduce the matrix scale function:
$\begin{split}S:\;&\mathcal{S}_{m}^{\mathbb{R}}\rightarrow\mathbb{R}^{+}\\ &\bm{\Sigma}\mapsto S(\bm{\Sigma})=s\end{split}$ (36)
satisfying the following assumptions [9, 10]:
  • A1
    Homogeneity of order one: $S(c\cdot\bm{\Sigma})=c\cdot S(\bm{\Sigma}),\forall c>0$,
  • A2
    Differentiability over $\mathcal{S}_{m}^{\mathbb{R}}$ with $\frac{\partial S(\bm{\Sigma})}{\partial[\bm{\Sigma}]_{11}}\neq 0$,
  • A3
    $S(\mathbf{I}_{N})=1$.
Then we define the shape matrix $\mathbf{V}_{S}$ as:
$\mathbf{V}_{S}\triangleq\bm{\Sigma}/S(\bm{\Sigma})\in\mathcal{S}_{m,S}^{\mathbb{R}},$ (37)
where $\mathcal{S}_{m,S}^{\mathbb{R}}$ is a (non-linear) differentiable manifold on dimension $m(m+1)/2-1$ such that:
$\mathcal{S}_{m,S}^{\mathbb{R}}\triangleq\{\mathbf{V}_{S}\in\mathcal{S}_{m}^{\mathbb{R}}|S(\mathbf{V}_{S})=1\}.$ (38)
We note in passing that the most popular choices for the scale function $S(\cdot)$ are $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. It is worth emphasizing that, under A1, A2 and A3, and as a direct consequence of the implicit function theorem, the first top-left entry of $\mathbf{V}_{S}$, i.e. $[\mathbf{V}_{S}]_{11}$, can (locally at $\mathbf{V}_{S}$) be expressed as function of the other entries. For example, if we choose $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, then $[\mathbf{V}_{S,0}]_{11}$ is trivially given by $[\mathbf{V}_{S,0}]_{11}=1$. For $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, we have that $[\mathbf{V}_{S}]_{11}=m-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{ii}$. Lastly, for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ it can be easily shown that, using the Laplace’s expansion of the determinant along the first row of $\mathbf{V}_{S}$, $[\mathbf{V}_{S}]_{11}$ can be recovered as $[\mathbf{V}_{S}]_{11}=\frac{1}{C_{11}}\left(1-\sum\nolimits_{i=2}^{m}[\mathbf{V}_{S}]_{1i}C_{1i}\right)$ where $C_{ij}$ indicates the cofactor of $[\mathbf{V}_{S}]_{ij}$.
As a consequence, the RES model can be parameterized in two different, yet equivalent, ways:
$\mathcal{P}_{\bm{\nu},g}=\left\{p_{X}(\mathbf{x}|\bm{\nu},g)=|\bm{\Sigma}|^{-1/2}g\left((\mathbf{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\nu}\in\Omega,g\in\overline{\mathcal{G}}\right\},$ (39)
and
$\mathcal{P}_{\bm{\eta},g}=\left\{p_{X}(\mathbf{x}|\bm{\eta},g)=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}g\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right);\bm{\eta}\in\Phi,g\in\overline{\mathcal{G}}\right\}$ (40)
where the finite-dimensional parameter of the first parameterization (39) makes use of the scatter matrix:
$\bm{\nu}\triangleq(\bm{\mu}^{T},\mathrm{vecs}(\bm{\Sigma})^{T})^{T}\in\Omega\subseteq\mathbb{R}^{m}\times\mathrm{vecs}(\mathcal{S}_{m}^{\mathbb{R}}),$ (41)
while the one of the second parameterization (40) is based on the shape matrix and on the scale:
$\bm{\eta}\triangleq(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T},s)^{T}\in\Phi\subseteq\mathbb{R}^{m}\times\underline{\mathrm{vecs}}(\mathcal{S}_{m,S}^{\mathbb{R}})\times\mathbb{R}^{+}.$ (42)
In both of the parameterizations, the infinite-dimensional parameter is the density generator $g\in\overline{\mathcal{G}}$.
It is important to note that the parameterization in (40) in terms of $(\underline{\mathrm{vecs}}(\mathbf{V}_{S}),s)$ is of interest for both practical and theoretical reasons. From a practical point of view, in a large part of applications such as Principal Component Analysis (PCA), Canonical Correlation Analysis (CCA) and subspace-based methods, only a scaled version of the covariance matrix, i.e. $\mathbf{V}_{S}$, is of interest while the scale term $s$ can be considered as a nuisance parameter. On the other hand, from a theoretical viewpoint, the parameterization in (40) allows us to investigate the hidden and counter-intuitive relationships between the parametric and semiparametric efficiency on the shape matrix $\mathbf{V}_{S}$ and the scale $s$ as in the analysis reported in [9, 10]. For this reason, in the following we will refer to $\mathcal{P}_{\bm{\eta},g}$ as the canonical parameterization of the RES model.
III-B The model $\mathcal{P}_{\bm{\eta},g}$ and its parametric and semiparametric submodels
The main reason for adopting the $\mathcal{P}_{\bm{\eta},g}$ model as canonical is that it can be analyzed using the geometric tools introduced in the Sect. II. In particular, in analogy to the models introduced in (6), (7), (8) and (9), and by indicating as
$\bm{\gamma}=(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T}\in\Gamma\subseteq\mathbb{R}^{m}\times\underline{\mathrm{vecs}}(\mathcal{S}_{m,S}^{\mathbb{R}})$ (43)
as the finite dimensional parameter vector of interest, we can define the following parametric submodel of $\mathcal{P}_{\bm{\eta},g}$:
$\mathcal{P}_{\bm{\gamma}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s_{0},g_{0}):\bm{\gamma}\in\Gamma\right\},$ (44)
$\mathcal{P}_{s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s,g_{0}):s\in\mathbb{R}^{+}\right\},$ (45)
$\mathcal{P}_{\bm{\gamma},s}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,g_{0}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+}\right\},$ (46)
along with the non-parametric model
$\mathcal{P}_{g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g):g\in\bar{\mathcal{G}}\right\}.$ (47)
where the scale $s\in\mathbb{R}^{+}$ and the density generator $g\in\bar{\mathcal{G}}$ play the role of finite and infinite-dimensional nuisance parameters respectively.
In the case of RES distributions, we can define an additional parametric model that is largely exploited in applications. In fact, since the density generator $g\in\bar{\mathcal{G}}$ is generally parameterized by a set of $p$ parameters, say $\bm{\zeta}\in\Delta\subseteq\mathbb{R}^{p}$ (for example the degrees of freedom of the $t$-distribution or the shape and scale parameters of the Generalized Gaussian distribution), one can suppose to know a priori the functional form of $g\in\bar{\mathcal{G}}$ up to its parameter vector $\bm{\zeta}\in\Delta$. This leads to the definition of the following parametric model:
$\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}=\left\{p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta}):\bm{\gamma}\in\Gamma,s\in\mathbb{R}^{+},\bm{\zeta}\in\Delta\right\}$ (48)
where
$p_{X}(\mathbf{x}|\bm{\gamma},s,\bm{\zeta})=s^{-m/2}|\mathbf{V}_{S}|^{-1/2}g_{\bm{\zeta}}\left(s^{-1}(\mathbf{x}-\bm{\mu})^{T}\mathbf{V}_{S}^{-1}(\mathbf{x}-\bm{\mu})\right)$ (49)
and where $\bm{\zeta}\in\Delta$ represents a second finite-dimensional nuisance term. This framework will allow us to apply Lemmas II.1 and II.2 to bring to light the counterintuitive links (in terms of efficiency) between the scale parameter $s$, the parameters $\bm{\zeta}$ of $g$ and $g$ itself. First, however, we must derive the three score vectors of the finite-dimensional parameters, that is $\mathbf{s}_{\bm{\gamma}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T})^{T}$, $s_{s_{0}}$ and $s_{\bm{\zeta}_{0}}$.
III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
Calculating the score vectors $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ and $s_{s_{0}}$ directly in $\mathcal{P}_{\bm{\eta},g}$ is not an easy task due to the fact that the differentiable manifold $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38) is not a linear subspace of the set of real matrices $\mathbb{R}^{m\times m}$. On the contrary, since $\mathcal{S}_{m}^{\mathbb{R}}$ is indeed a linear subspace of $\mathbb{R}^{m\times m}$, the score vector $\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}$ for the scatter matrix $\bm{\Sigma}$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) is easy to derive and well-known in the literature (see e.g. [22]). Fortunately, due to the differentability (under A1, A2 and A3) of $\mathcal{S}_{m,S}^{\mathbb{R}}$, we can define a diffeomorphism, say $\mathbf{w}$, between the canonical parameterisation of $\mathcal{P}_{\bm{\eta},g}$ and the one of $\mathcal{P}_{\bm{\nu},g}$. This allows us to easily pass from the score vectors of $\mathcal{P}_{\bm{\nu},g}$ to those of $\mathcal{P}_{\bm{\eta},g}$. Let us start by providing an explicit expression for such diffeomorphism.
By recalling that $\bm{\Sigma}=s\mathbf{V}_{S}$, under A1, A2 and A3, we can move from the parameterization (40) to the one in (39) (and vice versa) by means of
$\begin{split}\mathbf{w}:\;&\Phi\rightarrow\Omega\\ &\bm{\eta}=(\bm{\mu}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T},s)^{T}\mapsto\mathbf{w}(\bm{\eta})=(\bm{\mu}^{T},s\cdot[\mathbf{V}_{S}]_{11},s\cdot\underline{\mathrm{vecs}}(\mathbf{V}_{S})^{T})^{T},\end{split}$ (50)
whose Jacobian matrix $\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})$, evaluated at $\bm{\eta}_{0}\in\Phi$, is given by:
$\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})=\left(\begin{array}[]{ccc}\mathbf{I}_{m}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&s_{0}\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S,0}]_{11}&[\mathbf{V}_{S,0}]_{11}\\ \mathbf{0}&s_{0}\mathbf{I}_{m(m+1)/2-1}&\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})\\ \end{array}\right)\triangleq\left(\begin{array}[]{cc}\mathbf{I}_{m}&\mathbf{0}\\ \mathbf{0}&\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})\\ \end{array}\right)$ (51)
where $s_{0}=S(\bm{\Sigma}_{0})$ and $\mathbf{V}_{S,0}=\bm{\Sigma}_{0}/s_{0}$. Furthermore, the 2nd diagonal component in (51), as shown in [10, Sect. 4] and, by using our notation in the Appendix 156, of this work is given by
$\begin{split}\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}[\mathbf{V}_{S,0}]_{11}=-\left.\frac{\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))}{\partial S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))/\partial[\mathbf{V}_{S}]_{11}}\right|_{\mathbf{V}_{S}=\mathbf{V}_{S,0}},\end{split}$ (52)
where $S([\mathbf{V}_{S}]_{11},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))$ denotes the scale function applied to $\mathbf{V}_{S}$ reconstructed from $[\mathbf{V}_{S}]_{11}$ and $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$. For further reference, let us recast $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$ as:
$\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})=\left[s_{0}\mathbf{K}_{\mathbf{V}_{S,0}}\;\mathrm{vecs}(\mathbf{V}_{S,0})\right],$ (53)
where $\mathbf{K}_{\mathbf{V}_{S}}$ is a block-matrix defined as:
$\mathbf{K}_{\mathbf{V}_{S}}\triangleq\left[\begin{array}[]{c}\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}\\ \mathbf{I}_{m(m+1)/2-1}\end{array}\right],$ (54)
evaluated at $\mathbf{V}_{S,0}$. As an example, we show in Appendix 162 that the term $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$ can be explicitly obtained for the three above-mentioned scale functions as:
  • $\mathbf{0}^{T}_{m(m+1)/2-1}$ for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$,
  • $-\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $-\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m})_{1}}$ for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.
The derivation of the gradient $\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})$ in (51) can be also found in [10, 9] as function of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S}}$ that can be linked to our results through the relation:
$\mathbf{M}_{S}^{\mathbf{V}_{S,0}}=\mathbf{K}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}.$ (55)
Due to the crucial importance of this matrix in the following sections, we collect in the Appendix .6 both its known properties (from [10, 9]) and some original ones, which will be extensively used in the rest of this paper.
Let us now introduce the score vector for $\bm{\nu}_{0}\in\Omega$ in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) as $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$. By applying the rules of the differential matrix calculus detailed in [23, Ch. 8], the Kronecker product and the vec operator [23, Ch. 3], and the stochastic representation theorem (31), we get the following expressions of the scores (see [22, eqs. (34) and (37)]):
$\mathbf{s}_{\bm{\mu}_{0}}=_{d}\sqrt{\mathcal{Q}}\varphi_{0}(\mathcal{Q})\bm{\Sigma}^{-1/2}_{0}\mathbf{u},$ (56)
$\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}=_{d}2^{-1}\mathbf{D}_{m}^{T}\left[\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2}\right]\left[\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-\mathrm{vec}(\mathbf{I}_{m})\right],$ (57)
where, following the notation adopted in [8], we defined:  [^2] To avoid confusion, it is important to note that the function $\varphi_{0}(t)$ is linked to the equivalent function $\psi_{0}(t)$ used in [22, 24] by the constant $-2$: $\varphi_{0}(t)=-2\psi_{0}(t)$.
$\varphi_{0}(t)\triangleq\frac{-2}{g_{0}(t)}\frac{dg_{0}(t)}{dt}.$ (58)
By using the diffeomorphism in (50) and its Jacobian in (51), the score vector for $\bm{\eta}_{0}\triangleq(\bm{\gamma}_{0}^{T},s_{0})^{T}\in\Phi$ in $\mathcal{P}_{\bm{\eta},g}$ can be obtained from the derived score vector $\mathbf{s}_{\bm{\nu}_{0}}$ in $\mathcal{P}_{\bm{\nu},g}$ as:
$\mathbf{s}_{\bm{\eta}_{0}}=[\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})]^{T}\mathbf{s}_{\bm{\nu}_{0}}=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ \left[\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})\right]^{T}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\end{array}\right).$ (59)
From the block-diagonal structure $\mathbf{J}[\mathbf{w}](\bm{\eta}_{0})$ in (51), we immediately have that the score vector $\mathbf{s}_{\bm{\mu}_{0}}$ for the location $\bm{\mu}_{0}$ is equal to (56). Moreover, from the block structure of $\mathbf{J}[\mathbf{w}](\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0})$, by the definition of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ in (55), and by using the fact that:
$\mathrm{vecs}(\mathbf{V}_{S,0})^{T}\mathbf{D}_{m}^{T}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathbf{A})=\mathrm{tr}(\mathbf{A})$ (60)
for any $m\times m$ symmetric matrix $\mathbf{A}$, we have that:
$\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-\mathrm{vec}(\mathbf{I}_{m})\right)$ (61)
and
$s_{s_{0}}=_{d}(2s_{0})^{-1}\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right).$ (62)
Now that we have the explicit expression for the score vector $\mathbf{s}_{\bm{\eta}_{0}}$ in the canonical model $\mathcal{P}_{\bm{\eta},g}$ we can easily evaluate the related FIM $\mathbf{I}_{\bm{\eta}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\eta}_{0}}\mathbf{s}_{\bm{\eta}_{0}}^{T}\}$. Before computing FIM, we define two functionals of the true density generator $g_{0}$:
$\alpha(g_{0})\triangleq\frac{E\{\mathcal{Q}^{2}\varphi^{2}_{0}(\mathcal{Q})\}}{m(m+2)},$ (63)
$\beta(g_{0})\triangleq\frac{E\{\mathcal{Q}\varphi^{2}_{0}(\mathcal{Q})\}}{m}.$ (64)
Then, from standard calculations and by using the independence between $\mathcal{Q}$ and $\mathbf{u}$ (along with the properties of $\mathbf{u}$), the FIM for $\bm{\eta}_{0}\in\Phi$ is given by:
$\mathbf{I}_{\bm{\eta}_{0}}=\left(\begin{array}[]{cc}\mathbf{I}_{\bm{\gamma}_{0}}&\mathbf{I}_{\bm{\gamma}_{0},s_{0}}\\ \mathbf{I}_{\bm{\gamma}_{0},s_{0}}^{T}&I_{s_{0}}\end{array}\right)=\left(\begin{array}[]{ccc}\mathbf{I}_{\bm{\mu}_{0}}&\mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}&\mathbf{I}_{\bm{\mu}_{0},s_{0}}\\ \mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}&\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}&\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\\ \mathbf{I}_{\bm{\mu}_{0},s_{0}}^{T}&\mathbf{I}^{T}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}&I_{s_{0}}\end{array}\right),$ (65)
[…]
$\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\}=s_{0}^{-1}\beta(g_{0})\mathbf{V}_{S,0}^{-1},$ (66)
$\mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}\}=\mathbf{0},$ (67)
$\mathbf{I}_{\bm{\mu}_{0},s_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}s_{s_{0}}\}=\mathbf{0},$ (68)
$\begin{split}\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}&\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}\}=\\ &=\frac{1}{4}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\alpha(g_{0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})+\left(\alpha(g_{0})-1\right)\mathrm{vec}(\mathbf{I}_{m})\mathrm{vec}(\mathbf{I}_{m})^{T}\right]\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\right]^{T},\end{split}$ (69)
$\begin{split}I_{s_{0}}\triangleq E_{0}\{(s_{s_{0}})^{2}\}=\frac{m(m+2)\alpha(g_{0})-m^{2}}{4s_{0}^{2}},\end{split}$ (70)
$\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\triangleq E_{0}\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}s_{s_{0}}\}=\frac{(m+2)\alpha(g_{0})-m}{4s_{0}}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1}).$ (71)
We note that this result is in line with the expression already provided in [10, eqs. (6) and (7)].
We conclude this section by noticing that the score vector $\mathbf{s}_{\bm{\zeta}_{0}}$ for the nuisance parameters $\bm{\zeta}_{0}\in\Delta$, characterizing de density generator $g_{\bm{\zeta}}$ in $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ in (48), can be straightforwardly obtained from (49) as:
$\mathbf{s}_{\bm{\zeta}_{0}}=_{d}\frac{1}{g_{\bm{\zeta}_{0}}(\mathcal{Q})}\bm{\nabla}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q}).$ (72)
Now that we have clearly defined the statistical models and sub-models to be analyzed along with their corresponding score vectors, we can move on to analyze their geometric relationships and the consequent impact in terms of statistical efficiency.
IV The geometry of $\mathcal{P}_{\bm{\eta},g}$ and some results on parametric and semiparametric information matrices
In this section, we will specialise Lemmas II.1 and II.2, introduced in Sect. II, to the study of the elliptical distributions. The semiparametric model from which we start is the “canonical” one, i.e. $\mathcal{P}_{\bm{\eta},g}$ given in (40), corresponding to (c.t.) $\mathcal{P}$ in (5), along with its parametric submodels $\mathcal{P}_{\bm{\gamma}}$ in (44) (c.t. $\mathcal{P}_{1}$ in (6)), $\mathcal{P}_{s}$ in (45) (c.t. $\mathcal{P}_{2}$ in (7)), $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) and the non-parametric model $\mathcal{P}_{g}$ in (47) (c.t. $\mathcal{P}_{3}$ in (9)). Moreover, we will investigate the link of the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$, introduced in (48), with the above-mentioned models. The key result that allows us to study the impact of finite and infinite dimensional nuisance parameters on the semiparametric efficiency in $\mathcal{P}_{\bm{\eta},g}$ for the parameter vector of interest $\bm{\gamma}_{0}$ is Lemma II.2. To specialize this lemma to our context, we need to draw the following connections:
  • the finite-dimensional vector of the parameters of interest is $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ and the related score
    $\mathbf{s}_{\bm{\gamma}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T})^{T},$ (73)
    where $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ are given in (56) and (61), respectively.
  • the finite-dimensional nuisance parameter is $s_{0}$ and the related score $s_{s_{0}}$ is the one given in (62). As a consequence, from (14), the orthogonal projection onto the finite dimensional nuisance tangent space $\mathcal{H}\supseteq\mathcal{T}_{s_{0}}=\{a\cdot s_{s_{0}},a\in\mathbb{R}\}$ (c.t. $\mathcal{T}_{2}$ in (12)) is given by:
    $\Pi(h|\mathcal{T}_{s_{0}})=E_{0}\{hs_{s_{0}}\}I_{s_{0}}^{-1}s_{s_{0}},\;h\in\mathcal{H},$ (74)
    where $I_{s_{0}}$ has been derived in (70).
  • The infinite-dimensional nuisance tangent $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$) is given in Appendix .7 as:
    $\begin{split}\mathcal{T}_{g_{0}}&=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(s_{0}^{-1}(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{V}_{S,0}^{-1}(\mathbf{x}-\bm{\mu}_{0})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})),a.s.\;\mathbf{x}\in\mathcal{X}\right\}\\ &=\left\{h\in\mathcal{H}|h\;\text{is $\sigma(\mathcal{Q})$-measurable}\right\},\end{split}$ (75)
    where we used the definition of the transformation $Q_{\bm{\mu},\bm{\Sigma}}(\mathbf{x})$ in (32) and where $\sigma(\mathcal{Q})\subset\mathfrak{B}(\mathcal{X})$ is the sub-$\sigma$-algebra generated by the random variable $\mathcal{Q}$ in (32). Moreover, from e.g. [25, Ch. 23, Def. 4], we have that the orthogonal projection of a generic element $h\in\mathcal{H}$ onto $\mathcal{T}_{g_{0}}$ can be obtained as
    $\Pi(h|\mathcal{T}_{g_{0}})=E\{h|\mathcal{Q}\},\;\forall h\in\mathcal{H}.$ (76)
    Consequently, $\mathcal{H}\ni h\perp\mathcal{T}_{g_{0}}$ iff $E\{h|\mathcal{Q}\}=0$.
We are now ready to prove the following (surprising) result:
Proposition IV.1.
The efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\eta},g}$ in (40) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$ in (46). Moreover, if the scale functional $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then the parametric FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ in the parametric model $\mathcal{P}_{\bm{\gamma}}$ in (44) is equal to the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for $\mathcal{P}_{\bm{\eta},g}$.
This surprising result had already been discovered in [10, 9] using a semiparametric generalization of Le Cam’s theory [20]. The geometric approach, based solely on the tools discussed in Sect. II, provides an alternative proof that may be more accessible to a wider audience, as it requires only a basic background in Hilbert spaces.
Proof:
From Lemma II.2, to prove the first part of Proposition IV.1, we need to show that
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}_{g_{0}},$ (77)
which, by (73) and (76), is equivalent verifying that:
$E\{[\mathbf{s}_{\bm{\mu}_{0}}-\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{s_{0}})]|\mathcal{Q}\}=\mathbf{0}.$ (78)
$E\{[\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})]|\mathcal{Q}\}=\mathbf{0},$ (79)
where the score vectors $\mathbf{s}_{\bm{\mu}_{0}}$ and $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ are given in (56) and (61), respectively. Let us start with the condition (78) on the location parameter $\bm{\mu}_{0}$. From the definition of the projection operator onto $\mathcal{T}_{s_{0}}$, i.e. $\Pi(\cdot|\mathcal{T}_{s_{0}})$, given in (74) and from the fact that $\mathbf{I}_{\bm{\mu}_{0},s_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\mu}_{0}}s_{s_{0}}\}=\mathbf{0}$ as indicated in (68), we immediately have that:
$\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{s_{0}})=\mathbf{I}_{\bm{\mu}_{0},s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=\mathbf{0}.$ (80)
Moreover, from the expression of the score $\mathbf{s}_{\bm{\mu}_{0}}$ given by (56), the independence between $\mathcal{Q}$ and $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and the property $E\{\mathbf{u}\}=\mathbf{0}$, we have:
$E\{\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{Q}\}=\mathbf{0}.$ (81)
It now follows from equations (80) and (81) that condition (78) is satisfied. then, the Condition (78) is satisfied.
Let us now focus on the condition (79). We start by calculating the projection of $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ onto the finite-dimensional tangent space $\mathcal{T}_{s_{0}}$. From (74), (71), (70) and (81), we get:
$\begin{split}\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})&=\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}I_{s_{0}}^{-1}s_{s_{0}}=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)=E\{\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})|\mathcal{Q}\}.\end{split}$ (82)
Moreover, from the expression of the score $\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ given by (61) and by using the independence between $\mathcal{Q}$ and $\mathbf{u}$ and $E\{\mathbf{u}\mathbf{u}^{T}\}=m^{-1}\mathbf{I}_{m}$, we have:
$\begin{split}E\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{Q}\}&=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(E\{\mathbf{u}\mathbf{u}^{T}|\mathcal{Q}\})-\mathrm{vec}(\mathbf{I}_{m})\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathbf{I}_{m})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)\\ &=E\{\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})|\mathcal{Q}\}.\end{split}$ (83)
Consequently, the condition in (79) follows. Hence, the condition of Lemma II.2, equivalent to (77) in this setting, is verified.
To prove the second part of Proposition IV.1, we need to show that, if $S(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$, then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. From (82), we have that:
$\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}\Leftrightarrow\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})=\mathbf{0}$ (84)
or equivalently $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$ iff $\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\in\text{Ker}\;\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\mathbf{V}_{S,0}}),a\in\mathbb{R}\}$. As already noted in [10, Th. 3.1], this is a direct consequence of the Property P4 of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ (see appendix .6). In fact, for $S\equiv S_{d}$, we have that $\mathbf{D}_{S_{d}}^{\bm{\Sigma}_{0}}=m^{-1}|\bm{\Sigma}_{0}|^{1/m}\bm{\Sigma}_{0}^{-1}=m^{-1}\mathbf{V}_{S,0}^{-1}$. Consequently, we immediately have that $\mathrm{vec}(\mathbf{V}_{S_{d},0}^{-1})\in\text{Ker}\;\mathbf{M}_{S_{d}}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ and then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. ∎
As a side result of this proof and of the fact that $\mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=\mathbf{0}$ and $\mathbf{I}_{\bm{\mu}_{0},s_{0}}=\mathbf{0}$, it is immediate to verify that knowing or not knowing the location vector $\bm{\mu}_{0}$ has no impact on the efficiency losses when we do parametric or semiparametric inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$. Moreover, since $\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{g_{0}})=\mathbf{0}$, the lack of knowledge of the density generator does not induce any loss of efficiency w.r.t. $\bm{\mu}_{0}$ (see e.g. the discussion in [22]).
Roughly speaking Proposition IV.1 tells us that:
  • When performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the scale $s_{0}$ (while knowing the density generator $g_{0}$) or not knowing neither the scale $s_{0}$ nor the density generator $g_{0}$, leads to the same efficiency losses.
  • If the scale $S_{d}(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$ is adopted, when performing inference on $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S_{d}}^{\mathbb{R}}$, knowing or not knowing the scale $s_{0}$ and/or the density generator $g_{0}$, does not lead to any asymptotic efficiency loss w.r.t. the case in which $s_{0}$ and/or $g_{0}$ are perfectly known. For this reason, the scale $S_{d}(\bm{\Sigma}_{0})$ is called ”canonical” in [10].
Let us now focus our attention on the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ in (48) in which we have two finite-dimensional nuisance parameters: the scale $s_{0}$ and the parameters of the density generator $\bm{\zeta}_{0}\in\Delta$. The difference between $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ and the semiparametric model $\mathcal{P}_{\bm{\eta},g}$ in (40) is in the fact that in $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ the functional form of the density generator $g$ is assumed to be known up to a finite-dimensional vector of parameters while $g$ is fully unknown in $\mathcal{P}_{\bm{\eta},g}$.
Proposition IV.2.
The efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\eta},g}$ in (40) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ in (48).
Proof:
Let us start by proving that the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$ in (46). To this end, we just need to apply the Lemma II.2 by substituting the infinite-dimensional nuisance tangent space $\mathcal{T}_{3}$ with the finite-dimensional one generated by the score vector $\mathbf{s}_{\bm{\zeta}_{0}}$ in (11), i.e.
$\mathcal{H}\supseteq\mathcal{T}_{\bm{\zeta}_{0}}=\mathrm{Span}\{[\mathbf{s}_{\bm{\zeta}_{0}}]_{1},\ldots,[\mathbf{s}_{\bm{\zeta}_{0}}]_{p}\}.$ (85)
Consequently, we need to show that:
$\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}_{\bm{\zeta}_{0}},$ (86)
or, equivalently, that:
$\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\zeta}_{0}})=E_{0}\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}\mathbf{I}_{\bm{\zeta}_{0}}^{-1}\mathbf{s}_{\bm{\zeta}_{0}}=\mathbf{0},$ (87)
where $\mathbf{I}_{\bm{\zeta}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\zeta}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}$ and $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$, from (82), can be expressed as:
$\begin{split}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}&=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ \mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})\end{array}\right)\\ &=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ 2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathcal{Q}\varphi_{0}(\mathcal{Q})[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-m^{-1}\mathrm{vec}(\mathbf{I}_{m})\right)\end{array}\right).\end{split}$ (88)
We can also introduce the efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ (c.t. (23)) in the parametric model $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance terms $s_{0}$ as:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})\triangleq E_{0}\left\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}^{T}\right\}=\left(\begin{array}[]{cc}\mathbf{I}_{\bm{\mu}_{0}}&\mathbf{0}\\ \mathbf{0}&\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\end{array}\right),$ (89)
whose block-diagonal structure follows once again from the properties of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and from the independence of $\mathbf{u}$ with $\mathcal{Q}$.
Clearly, the equality in (87) is verified iff $E_{0}\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}=\mathbf{0}$. This condition can be shown to hold using the mutual independence of $\mathcal{Q}$ and $\mathbf{u}$ and from the properties of $\mathbf{u}$. In fact, from the expression of $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ given in (88), we have that:
$\begin{split}E_{0}&\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}\\ &=\left(\begin{array}[]{c}\bm{\Sigma}^{-1/2}_{0}E\left\{\frac{\sqrt{\mathcal{Q}}\varphi_{0}(\mathcal{Q})}{g_{\bm{\zeta}_{0}}(\mathcal{Q})}\bm{\nabla}^{T}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q})\right\}E\{\mathbf{u}\}\\ 2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]E\left\{\frac{\mathcal{Q}\varphi_{0}(\mathcal{Q})}{g_{\bm{\zeta}_{0}}^{2}(\mathcal{Q})}\bm{\nabla}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q})\bm{\nabla}_{\bm{\zeta}}^{T}g_{\bm{\zeta}_{0}}(\mathcal{Q})\right\}\left(\mathrm{vec}(E\{\mathbf{u}\mathbf{u}^{T}\})-m^{-1}\mathrm{vec}(\mathbf{I}_{m})\right)\end{array}\right)=\mathbf{0}.\end{split}$ (90)
since $E\{\mathbf{u}\}=\mathbf{0}$ and $E\{\mathbf{u}\mathbf{u}^{T}\}=m^{-1}\mathbf{I}_{m}$. Finally, the statement of Proposition IV.2 follow immediately from the fact that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$, as shown in Proposition IV.1. ∎
Roughly speaking Proposition IV.2 tells us that, if the scale $s_{0}$ is an unknown parameter, when performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the parameters $\bm{\zeta}_{0}$ of the density generator $g_{\bm{\zeta}_{0}}$ or not knowing the whole functional form $g$, leads to the same efficiency losses.
Propositions IV.1 and IV.2 give us a clear picture of the efficiency relationships among the various sub-models of elliptical distributions in terms of Fisher information matrices. It would now be interesting to obtain explicit closed-form expressions for the inverses of these FIM matrices in order to obtain the related information bounds. This will be the objective of the next section.
V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$ and $\mathcal{P}_{\bm{\nu},g}$
The purpose of this section is to provide the counterpart of the Propositions IV.1 and IV.2 in terms of information bounds.
A clarification is in order before continuing. In the theory of parametric estimation, the best known information bound is the Cramér-Rao Bound (CRB) which coincides with inverse of the related FIM as shown e.g. [17, Chap. 5, Sects 5 and 6]. Consequently, for the different parametric submodels of the canonical semiparametric model in $\mathcal{P}_{\bm{\eta},g}$ in (40), we can derive the related information/CR bounds as:
  • CRB for the parameters of interest $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ in the presence of the finite-dimensional nuisance $s_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\gamma},s}$ in (46):
    $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})^{-1},$ (91)
    where $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ is the efficient FIM as defined in (89).
  • CRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of two finite-dimensional nuisance terms, $s_{0}$ and $\bm{\zeta}_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ in (48):
    $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})^{-1}=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})^{-1}$ (92)
    as shown in Proposition IV.2.
When we move to the semiparametric case, the CRB can no longer be defined as in the classical parametric case. Remarkably, the Hájek-Le Cam convolution theorem (see e.g. [1, Sect. 3.3, Theo. 2]) provides the right theoretical framework to unify the concept of information bound in the parametric, semiparametric and non-parametric case. A formal presentation of this theorem would lead us too far from the main purpose of this article. Therefore, below we will simply define the “Semiparametric CRB (SCRB)” as the information bound obtained as the inverse of the efficient semiparametric FIM (SFIM) (for an in-depth discussion about this point, we refer the readers to [1, Chap. 3]). Specifically, for the semiparametric canonical model $\mathcal{P}_{\bm{\eta},g}$ in (40), we have that the SCRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of a finite-dimensional $s_{0}$ and of in infinite-dimensional $g_{0}$ nuisance terms is given by:
$\mathrm{SCRB}(\bm{\gamma}_{0}|s_{0},g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})^{-1},$ (93)
where $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ is the efficient Semiparametric FIM as defined in (22).
However, as said before, the canonical parameter of interest $\bm{\gamma}_{0}$ contains the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ that belongs to the differentiable but non-linear manifold $\mathcal{S}_{m,S}^{\mathbb{R}}$ in (38) and this causes considerable difficulties when conducting inference on $\mathbf{V}_{S,0}$. For this reason, it is often preferable to make a “constrained” inference in the model $\mathcal{P}_{\bm{\nu},g}$ in (39) where the parameter $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ contains the scatter matrix $\bm{\Sigma}_{0}$ belonging to $\mathcal{S}_{m}^{\mathbb{R}}$ that is indeed a linear subspace of $\mathbb{R}^{m\times m}$. Clearly, performing inference on $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$ in $\mathcal{P}_{\bm{\eta},g}$ in the presence of the nuisance parameters $s_{0}$ and $g_{0}$ is equivalent (at least locally at $\mathbf{V}_{S,0}$, thanks to the implicit function theorem) performing inference on ${\bm{\Sigma}}_{0}\in\mathcal{S}_{m}^{\mathbb{R}}$ in $\mathcal{P}_{\bm{\nu},g}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ satisfying the assumptions A1, A2 and A3 in Sect. III. This highlights the usefulness of the two constrained information/CR bounds:
  • Constrained parametric CRB (CCRB) for the parameters $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\nu},g_{0}}$. This bound is well-defined in the standard parametric theory as shown in [26].
  • Constrained Semiparametric CRB (CSCRB) for the parameters $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ in the semiparametric submodel $\mathcal{P}_{\bm{\nu},g}$ in the presence of the infinite-dimensional nuisance $g_{0}\in\bar{\mathcal{G}}$. This bound is well-defined by a constrained generalization of the Hájek-Le Cam convolution theorem as shown in [27].
The following three subsections are organized as follows: in subsect. V-A, an explicit expression for the parametric CRB of $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ in the presence of the nuisance parameter $s_{0}$ is provided, while in subsect. V-B the constrained parametric CRB for $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}$ under the constraint $S(\bm{\Sigma}_{0})=s_{0}$ is given. Finally, the subsect. V-C proposes the counterpart of the Propositions IV.1 and IV.2 in terms of parametric and semiparametric information bounds.
V-A Explicit expression for $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$
To calculate (in a closed-form not containing matrix inverses) the $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in (91), we will exploit the expression of the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma})_{0}^{T})^{T}$ deduced from the inverse of the FIM $\mathbf{I}(\bm{\nu}_{0})$. To this end, let us consider the inverse, say $\mathbf{w}^{-1}$, of the diffeomorphism $\mathbf{w}$ given in (50). It can be explicitly obtained as:
$\begin{split}\mathbf{w}^{-1}:\;&\Omega\rightarrow\Phi\\ &\bm{\nu}=(\bm{\mu}^{T},\mathrm{vecs}(\bm{\Sigma})^{T})^{T}\mapsto\mathbf{w}^{-1}(\bm{\nu})=(\bm{\mu}^{T},S^{-1}(\bm{\Sigma})\underline{\mathrm{vecs}}(\bm{\Sigma})^{T},S(\bm{\Sigma}))^{T},\end{split}$ (94)
whose Jacobian matrix $\mathbf{J}[\mathbf{w}^{-1}](\bm{\nu}_{0})$ is given by:
$\mathbf{J}[\mathbf{w}^{-1}](\bm{\nu}_{0})=\left(\begin{array}[]{cc}\mathbf{I}_{m}&\mathbf{0}\\ \mathbf{0}&S^{-1}(\bm{\Sigma}_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\left[\mathbf{I}_{m^{2}}-\mathrm{vec}(\mathbf{V}_{S,0})\bm{\nabla}_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma}_{0})\right]\mathbf{D}_{m}\\ \mathbf{0}&\bm{\nabla}_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma}_{0})\mathbf{D}_{m}\\ \end{array}\right).$ (95)
The following proposition is proved in Appendix 199:
Proposition V.1.
The parametric CRB for $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ is given by
$\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})\triangleq\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})^{-1}=\left(\begin{array}[]{cc}\mathrm{CRB}(\bm{\mu}_{0})&\mathbf{0}\\ \mathbf{0}&\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\\ \end{array}\right),$ (96)
where:
$\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\frac{s_{0}}{\beta(g_{0})}\mathbf{V}_{S,0}$ (97)
and
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (98)
with $\mathbf{P}_{S}(\mathbf{V}_{S,0})\triangleq\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0})\bm{\nabla}_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})$, which takes the following expressions:
  • $\mathbf{I}_{m^{2}}-{\rm vec}(\mathbf{V}_{S,0}){\bf e}^{T}_{1}$ for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{I}_{m^{2}})^{T}$ for $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$,
  • $\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0}^{-1})^{T}$ for $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$.
Furthermore, for the scale function $S_{d}(\bm{\Sigma})\triangleq|\bm{\Sigma}|^{1/m}$, the parameters $\mathbf{V}_{S,0}$ and $s_{0}$ are decoupled in the CRB for the parameter pair $(\mathbf{V}_{S,0},s_{0})$ and the matrix $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ reduces to
$\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0})|s_{d,0})=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\left[(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})-2m^{-1}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0})^{T}\right]\mathbf{D}_{m}^{\#^{T}}\underline{\mathbf{I}}_{m}^{T}.$ (99)
We note that the proof of Proposition V.1 given in Appendix 199 also provides us with closed-form expressions of CRB$(\mathrm{vecs}(\bm{\Sigma}_{0}))$ (187) and $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))$ (197) (199), which are new results.
V-B The Constrained parametric CRB (CCRB) in $\mathcal{P}_{\bm{\nu},g_{0}}$ under $S(\bm{\Sigma}_{0})=s_{0}$
As shown in Appendix 199, the FIM for $\bm{\nu}_{0}\in\Omega$ in the parametric model $\mathcal{P}_{\bm{\nu},g_{0}}$ in (39), i.e. $\mathbf{I}_{\bm{\nu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\nu}_{0}}\mathbf{s}_{\bm{\nu}_{0}}^{T}\}$ in (178) is block-diagonal. Consequently the CCRB on $\bm{\nu}_{0}\in\Omega$ has, in turn, the following block-diagonal structure:
$\mathrm{CCRB}(\bm{\nu}_{0})=\left(\begin{array}[]{cc}\mathrm{CRB}(\bm{\mu}_{0})&\mathbf{0}\\ \mathbf{0}&\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\\ \end{array}\right),\;S(\bm{\Sigma}_{0})=s_{0},$ (100)
where $\mathrm{CRB}(\bm{\mu}_{0})$ is the one already given in (97).
To derive $\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))$, let’s take a step back to the term $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$. In fact, as an alternative method w.r.t. the one used in (98), $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ can be derived by exploiting the block-structure of the FIM $\mathbf{I}_{\bm{\eta}_{0}}$ in (65). Specifically, from (65)-(71) and by recalling the definition of the efficient FIM $\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ in (89), we have:
$\begin{split}\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})&\triangleq\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}=\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-I^{-1}_{s_{0}}\mathbf{I}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\mathbf{I}^{T}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}),s_{0}}\\ &=2^{-1}\alpha(g_{0})\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\frac{(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})}{2}-\frac{1}{m}\mathrm{vec}(\mathbf{I}_{m})\mathrm{vec}(\mathbf{I}_{m})^{T}\right]\left[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}\right]\left[\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\right]^{T},\\ &=2^{-1}\alpha(g_{0})\mathbf{K}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{K}_{\mathbf{V}_{S,0}},\\ &=2^{-1}\alpha(g_{0})\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1},\end{split}$ (101)
where:
  1. 1.
    in the second identity, we used (55) and the two properties of the commutation matrix $\mathbf{K}_{m}$, i.e. i) $\mathbf{K}_{m}(\mathbf{A}\otimes\mathbf{B})=(\mathbf{B}\otimes\mathbf{A})\mathbf{K}_{m}$ and ii) $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$ [23, Ch. 3],
  2. 2.
    in the third identity, we introduced the matrix:
    $\mathbf{T}_{\mathbf{A}}\triangleq\mathbf{A}^{-1}\otimes\mathbf{A}^{-1}-m^{-1}\mathrm{vec}(\mathbf{A}^{-1})\mathrm{vec}(\mathbf{A}^{-1})^{T}$ (102)
    for any invertible $m\times m$ matrix $\mathbf{A}$
  3. 3.
    finally, in the forth identity, we used (165) defined in Appendix .6.
Consequently, from the basic properties of the CRB (see e.g. [17, Th. 6.6]) and from (164) in Appendix .6, we have that the CRB on $\mathrm{vecs}(\mathbf{V}_{S,0})$ can be obtained as:
$\begin{split}\mathrm{CRB}(\mathrm{vecs}(\mathbf{V}_{S,0})|s_{0})&=\mathbf{K}_{\mathbf{V}_{S,0}}\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\mathbf{K}_{\mathbf{V}_{S,0}}^{T}\\ &=\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\left[\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\right]^{-1}\mathbf{S}_{\mathbf{V}_{S,0}}^{-T}\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{S}_{\mathbf{V}_{S,0}}^{-1}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}\left[\mathbf{U}_{\mathbf{V}_{S,0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\mathbf{V}_{S,0}}\mathbf{D}_{m}\mathbf{U}_{\mathbf{V}_{S,0}}\right]^{-1}\mathbf{U}_{\mathbf{V}_{S,0}}^{T},\end{split}$ (103)
where we used (167) and the fact that $\mathbf{S}_{\mathbf{V}_{S,0}}$ is, by definition, invertible.
Let us now consider the constrained scatter matrix $\bm{\Sigma}_{0}=s_{0}\mathbf{V}_{S,0}$, such that $S(\bm{\Sigma}_{0})=s_{0}$. Using again the properties of the CRB [17, Theo. 6.6] and the fact that, as discussed in the property P5 in Appendix .6, $\mathbf{S}_{\mathbf{V}_{S,0}}=\mathbf{S}_{\bm{\Sigma}_{0}}$ and $\mathbf{U}_{\mathbf{V}_{S,0}}=\mathbf{U}_{\bm{\Sigma}_{0}}$ we have that the CCRB on the $\bm{\Sigma}_{0}$, under the constraint $S(\bm{\Sigma}_{0})=s_{0}$, is given by:
$\begin{split}\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))&=\left[s_{0}\mathbf{K}_{\mathbf{V}_{S,0}}\right]\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})^{-1}\left[s_{0}\mathbf{K}_{\mathbf{V}_{S,0}}\right]^{T}\\ &=2\alpha(g_{0})^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}\left[\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\bm{\Sigma}_{0}}\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}_{0}}\right]^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}^{T},\quad\text{for}\;S(\bm{\Sigma}_{0})=s_{0}.\end{split}$ (104)
If we compare (after carefully harmonising the notations) the $\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))$ for $S(\bm{\Sigma}_{0})=s_{0}$ obtained in (104) and the Constrained Semiparametric CRB on $\mathrm{vecs}(\bm{\Sigma}_{0})$ obtained in [22, eq. (59)] we can recognize that they are the same. This surprising result is a direct consequence of the intrinsic geometry of the semiparametric models $\mathcal{P}_{\bm{\eta},g}$ and $\mathcal{P}_{\bm{\nu},g}$ that has been brought to light by Propositions IV.1 and IV.2. To summarize and emphasize the results obtained in this section, let us conclude with the counterpart for the information bounds of Propositions IV.1 and IV.2.
V-C Three equality chains for parametric and semiparametric information bounds in RES distributions
We are finally ready to introduce three sequences of equalities among information bounds for the parameters of interest in RES distributions. The following proposition summarizes the key points from the previous sections and can serve as a take-away message for readers who are not interested in full mathematical details.
Proposition V.2.
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector with location vector $\bm{\mu}_{0}$, scatter matrix $\bm{\Sigma}_{0}$ and density generator $g_{0}\in\bar{\mathcal{G}}$. Let $S(\bm{\Sigma}_{0})=s_{0}$ be a constraint on the covariance matrix where $S$ is a matrix function satisfying Assumptions A1, A2 and A3 and let $\mathbf{V}_{S,0}=s_{0}^{-1}\bm{\Sigma}_{0}$ the related shape matrix. Moreover:
  • Let $\mathbf{D}_{S}^{\bm{\Sigma}_{0}}\triangleq\frac{\partial S(\bm{\Sigma}_{0})}{\partial\bm{\Sigma}_{0}}$,
  • Let $\mathbf{U}_{\bm{\Sigma}_{0}}$ be a matrix whose columns form an orthonormal basis for the null space of $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}_{0}})^{T}$, i.e.:
    $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}_{0}})^{T}\mathbf{U}_{\bm{\Sigma}_{0}}=\mathbf{0},\quad\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{U}_{\bm{\Sigma}_{0}}=\mathbf{I}_{m(m+1)/2-1}.$
  • $\mathbf{T}_{\bm{\Sigma}_{0}}\triangleq\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}-m^{-1}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}$,
Then, the following chain of equalities hold:
  1. 1.
    Relation among the SCRB for the location vector $\bm{\mu}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\nu},g}$ (or $\mathcal{P}_{\bm{\eta},g}$) and three different CRB in the parametric models $\mathcal{P}_{\bm{\nu},s}$ (or $\mathcal{P}_{\bm{\eta},s}$), $\mathcal{P}_{\bm{\nu},s,\bm{\zeta}}$ (or $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$) and $\mathcal{P}_{\bm{\nu}}$ (or $\mathcal{P}_{\bm{\eta}}$):
    $\begin{split}&\beta(g_{0})^{-1}\bm{\Sigma}_{0}\\ &=\mathrm{SCRB}(\bm{\mu}_{0}|s_{0},g_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally unknown}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}|s_{0},\bm{\zeta}_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}|s_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CRB}(\bm{\mu}_{0}),\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (105)
  2. 2.
    Relation among the constrained SCRB (CSCRB) for the constrained scatter matrix $\bm{\Sigma}_{0}$ in the semiparametric model $\mathcal{P}_{\bm{\nu},g}$ and three different constrained CRB (CCRB) in the parametric models $\mathcal{P}_{\bm{\nu},s}$, $\mathcal{P}_{\bm{\nu},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\nu}}$, under $S(\bm{\Sigma}_{0})=s_{0}$:
    $\begin{split}2\alpha(g_{0})^{-1}&\mathbf{U}_{\bm{\Sigma}_{0}}\left[\mathbf{U}_{\bm{\Sigma}_{0}}^{T}\mathbf{D}_{m}^{T}\mathbf{T}_{\bm{\Sigma}_{0}}\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}_{0}}\right]^{-1}\mathbf{U}_{\bm{\Sigma}_{0}}^{T},\quad\text{for}\;S(\bm{\Sigma}_{0})=s_{0}.\\ &=\mathrm{CSCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0},g_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionallyunknown}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0},\bm{\zeta}_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})|s_{0}),\;\forall S(\bm{\Sigma}_{0})=s_{0}\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CCRB}(\mathrm{vecs}(\bm{\Sigma}_{0})),\;\textit{iff}\;|\bm{\Sigma}_{0}|^{1/m}=s_{0}\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (106)
  3. 3.
    Relation among the SCRB for $\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})$ in the canonical semiparametric model $\mathcal{P}_{\bm{\eta},g}$ and three different parametric CRB (CRB) in the parametric models $\mathcal{P}_{\bm{\eta},s}$, $\mathcal{P}_{\bm{\eta},s,\bm{\zeta}}$ and $\mathcal{P}_{\bm{\eta}}$:
    $\begin{split}\alpha^{-1}(g_{0})&\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},\\ &=\mathrm{SCRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0},g_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally unknown}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0},\bm{\zeta}_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{functionally known up to its parameters}\;\bm{\zeta}_{0}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0}),\quad\left[s_{0}\;\text{unknown},\;g_{0}\;\text{fully known}\right]\\ &=\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0})),\;\textit{iff}\;|\mathbf{V}_{S_{d},0}|^{1/m}=1\quad\left[s_{0}\;\text{known},\;g_{0}\;\text{fully known}\right].\end{split}$ (107)
Proof:
The proof of the Proposition V.2 follow directly from the Propositions IV.1 and IV.2 and in particular from the fact that:
  1. 1.
    $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$,
  2. 2.
    If $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then $\bar{\mathbf{I}}(\bm{\gamma}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$.
The proof is concluded by noticing that the SCRB and the different CRBs are defined as the inverse of the related FIM according to the Hájek-Le Cam convolution theorem for parametric models [1, Sect. 2.3, Th. 1], semiparametric models [1, Sect. 3.3, Theo. 2] and their constrained counterparts [27]. ∎
VI Parameterization of the location vector and scatter matrix
In this section, we focus our attention to the case where both the location vector and the scatter matrix can be parametrized by a real $d$-dimensional parameter vector $\bm{\theta}=(\bm{\gamma}^{T},\bm{\xi}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi\subset\mathbb{R}^{d}$, where $d=q+r$.
VI-A Some preliminaries
Let us start with some preliminaries. Let $\mathcal{X}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g)$ be a RES-distributed random vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})\in\mathbb{R}^{m}$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})\in\mathcal{S}_{m}^{\mathbb{R}}$ are parameterized by a $d$-dimensional parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Theta\triangleq\Gamma\times\Psi$. As in Sect. II, $\Gamma\subseteq\mathbb{R}^{q}$ denotes the set of the (finite-dimensional) parameter vectors $\bm{\gamma}$ of interests, $\Psi\subseteq\mathbb{R}^{r}$ denotes the set of (finite-dimensional) nuisance parameter vectors $\bm{\xi}$ and $\bar{\mathcal{G}}$ is the set of the (infinite-dimensional) nuisance functions $g$ already defined in (34). In the following, we always assume that the parameterization $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ satisfies the following assumptions:
  • P1)
    it is continuous on $\Theta$,
  • P2)
    the two Jacobian matrices $\mathbf{J}[\bm{\mu}_{0}]\triangleq\mathbf{J}[\bm{\mu}(\bm{\theta}_{0})]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]\triangleq\mathbf{J}[\mathrm{vec}(\bm{\Sigma}(\bm{\theta}_{0}))]$ are full column rank. This ensures that the parameterisation is locally one-to-one in an open neighbourhood of $\bm{\theta}_{0}\in\Theta$.
  • P3)
    the inverse $[\bm{\Sigma}(\bm{\theta})]^{-1}$ exists for all $\bm{\theta}\in\Theta$.
The related semiparametric model can be then expressed as:
$\mathcal{P}_{\bm{\theta},g}=\left\{p_{X}(\mathbf{x}|,\bm{\theta},g)=|\bm{\Sigma}(\bm{\theta})|^{-1/2}g\left((\mathbf{x}-\bm{\mu}(\bm{\theta}))^{T}\bm{\Sigma}(\bm{\theta})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\theta}))\right);\bm{\theta}\in\Theta,g\in\overline{\mathcal{G}}\right\},$ (108)
where, from (32), we have that:
$(\mathbf{x}-\bm{\mu}(\bm{\theta}_{0}))^{T}\bm{\Sigma}(\bm{\theta}_{0})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\theta}_{0}))\triangleq Q_{0}=_{d}\mathcal{Q}.$ (109)
In line with our earlier analysis, we may pose the following question: is it possible to characterize all parameterisations $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ that imply that not knowing the finite-dimensional nuisance vector $\bm{\xi}_{0}\in\Psi$, while knowing $g_{0}$, leads to the same loss of efficiency as not knowing $\bm{\xi}_{0}\in\Psi$ and the density generator $g_{0}\in\overline{\mathcal{G}}$ (i.e. the infinite-dimensional nuisance parameter)?
To answer this question, let us start by evaluating the score vector
$\mathbf{s}_{\bm{\theta}_{0}}=(\mathbf{s}_{\bm{\gamma}_{0}}^{T},\;\mathbf{s}_{\bm{\xi}_{0}}^{T})^{T},$ (110)
and the related FIM $\mathbf{I}_{\bm{\theta}_{0}}={E_{0}\{\mathbf{s}_{\bm{\theta}_{0}}\mathbf{s}_{\bm{\theta}_{0}}^{T}\}}$ of the parametric model $\mathcal{P}_{\bm{\theta},g_{0}}$. From the definition of the parametric and non-parametric submodels of a semiparametric model given in Sect. II, it is immediate to verify that $\mathcal{P}_{\bm{\theta},g_{0}}$ correspond to the parametric submodel $\mathcal{P}_{1,2}$ in (8). Moreover, as a direct consequence of (109), we have that the non-parametric submodel $\mathcal{P}_{g}$ of $\mathcal{P}_{\bm{\theta},g}$ (that correspond to $\mathcal{P}_{3}$ in (9)) of $\mathcal{P}_{\bm{\theta},g}$ admits the same tangent space $\mathcal{T}_{g_{0}}$ already derived in (75). As the consequence, the orthogonal projection of a generic element $h\in\mathcal{H}$ onto $\mathcal{T}_{g_{0}}$ is again the conditional expectation given in (76).
Let us start with the calculation of the score function $\mathbf{s}_{\bm{\theta}_{0}}$. Following the derivation in [28, Sec. 3.1] and [29, Sec. III], each entry of $\mathbf{s}_{\bm{\theta}_{0}}$ can be easily evaluated as:
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}&\triangleq\left.\frac{\partial\ln p_{X}\left(\mathbf{x};\bm{\theta},g_{0}\right)}{\partial\theta_{i}}\right|_{\bm{\theta}=\bm{\theta}_{0}}=-2^{-1}\left(\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(Q_{0})Q_{i}^{0}\right),\end{split}$ (111)
where, according to the adopted notation $Q_{i}^{0}\triangleq\frac{\partial Q_{0}}{\partial\theta_{i}}$, the function $\varphi_{0}$ has already been defined in (58), and
$\mathbf{P}_{i}^{0}\triangleq\bm{\Sigma}_{0}^{-1/2}\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1/2}.$ (112)
Moreover, from direct calculation, we have:
$Q_{i}^{0}=-2(\mathbf{x}-\bm{\mu}_{0})^{T}\bm{\Sigma}_{0}^{-1}\bm{\mu}_{i}^{0}-(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{S}_{i}^{0}(\mathbf{x}-\bm{\mu}_{0}),$ (113)
where $\bm{\mu}_{i}^{0}\triangleq\frac{\partial\bm{\mu}_{0}}{\partial\theta_{i}}$ and
$\mathbf{S}_{i}^{0}=\bm{\Sigma}_{0}^{-1}\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1}.$ (114)
By collecting previous results, the entries of the score vector $\mathbf{s}_{\bm{\theta}_{0}}$ can be expressed as:
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}=-2^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(Q_{0})(\mathbf{x}-\bm{\mu}_{0})^{T}\bm{\Sigma}_{0}^{-1}\bm{\mu}_{i}^{0}+2^{-1}\varphi_{0}(Q_{0})(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{S}_{i}^{0}(\mathbf{x}-\bm{\mu}_{0}),\;i=1,\ldots,d.\end{split}$ (115)
Using the stochastic representation in (31), the score vector in (115) can be rewritten as:
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}=_{d}-2^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(\mathcal{Q})\left(\sqrt{\mathcal{Q}}\mathbf{u}^{T}\bm{\Sigma}_{0}^{-1/2}\bm{\mu}_{i}^{0}+2^{-1}\mathcal{Q}\mathbf{u}^{T}\mathbf{P}_{i}^{0}\mathbf{u}\right),\;i=1,\ldots,d.\end{split}$ (116)
Consequently, the entries of the $q$-dimensional score vector of the parameters of interest can be simply expressed as:
$\begin{split}[\mathbf{s}_{\bm{\gamma}_{0}}]_{i}=[\mathbf{s}_{\bm{\theta}_{0}}]_{i}\;i=1,\ldots,q,\end{split}$ (117)
while the $r$-dimensional score vector of the finite-dimensional nuisance parameters is given by:
$\begin{split}[\mathbf{s}_{\bm{\xi}_{0}}]_{j}=[\mathbf{s}_{\bm{\theta}_{0}}]_{q+j}\;j=1,\ldots,r.\end{split}$ (118)
To derive the FIM $\mathbf{I}_{\bm{\theta}_{0}}$, we may use the procedure in [29] that leads to the following compact expression given in [8, Sect. 1.6.5] as:
$\begin{split}\mathbf{I}_{\bm{\theta}_{0}}=\beta(g_{0})&\mathbf{J}[\bm{\mu}_{0}]^{T}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\\ &+2^{-1}\alpha(g_{0})\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}+2^{-1}(1-\alpha(g_{0})^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]\end{split}$ (119)
where the scalars $\beta(g_{0})$ and $\alpha(g_{0})$ are given in (64) and (63) respectively, while the Jacobian matrices, $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, are explicitly expressed as
$[\mathbf{J}[\bm{\mu}_{0}]]_{i,j}=\left.\frac{\partial[\bm{\mu}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}},\;[\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]]_{i,j}=\left.\frac{\partial[\mathrm{vec}(\bm{\Sigma}(\bm{\theta}))]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$ (120)
It is important to note that the assumption P2 of full column rank for $\mathbf{J}[\bm{\mu}_{0}]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]$ guarantees that $\mathbf{I}_{\bm{\theta}_{0}}$ is invertible.
VI-B Conditions of equality between efficient SFIM and parametric efficient FIM
By collecting the previous results, we are ready to state the following proposition whose proof is given in Appendix 203:
Proposition VI.1.
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})$ are parameterized by $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Gamma\times\Psi$, where $\bm{\gamma}_{0}\in\Gamma$ is the $q$-dimensional vector of interest and $\bm{\xi}_{0}\in\Psi$ is the $r$-dimensional nuisance vector, such that $d=q+r$. Let $g_{0}\in\overline{\mathcal{G}}$ be the infinite-dimensional nuisance parameter. Then, the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\theta},g}$ in (108) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in $\mathcal{P}_{\bm{\theta},g_{0}}$ iff the following condition is satisfied:
$\left(\mathbf{J}_{\bm{\gamma}}^{T}[\mathrm{vec}(\bm{\Sigma}_{0})]-\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}\mathbf{J}_{\bm{\xi}}^{T}[\mathrm{vec}(\bm{\Sigma}_{0})]\right)\mathrm{vec}(\bm{\Sigma}_{0}^{-1})=\mathbf{0},$ (121)
where $\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}$ and $\mathbf{I}_{\bm{\xi}_{0}}$ are two sub-blocks of the FIM $\mathbf{I}_{\bm{\theta}_{0}}=\left(\begin{array}[]{cc}\mathbf{I}_{\bm{\gamma}_{0}}&\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\\ \mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}^{T}&\mathbf{I}_{\bm{\xi}_{0}}\end{array}\right)$ given in (119).
As examples of how condition (121) can be used, let us consider two quite common parameterization in signal processing application: the case where the location vector and the scatter matrix have no parameters in common and the “low-rank”, parameterization model.
VI-C The elliptical parameterized model where the location vector and the scatter matrix have no parameters in common
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ is parameterized by the parameter of interest $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ while the scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$ is parameterized by the nuisance parameter $\bm{\xi}_{0}\in\Psi\subseteq\mathbb{R}^{r}$. Furthermore, we assume that such parameterisation satisfies the assumptions P1 and P2 given in Sect. VI-A.
Let us consider the three following models:
$\mathcal{P}_{\bm{\gamma},\bm{\xi},g}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g)=|\bm{\Sigma}(\bm{\xi})|^{-1/2}g((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,g\in\overline{\mathcal{G}}\right\},$ (122)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g_{0})=|\bm{\Sigma}(\bm{\xi})|^{-1/2}g_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\},$ (123)
$\mathcal{P}_{\bm{\gamma}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi}_{0},g_{0})=|\bm{\Sigma}(\bm{\xi}_{0})|^{-1/2}g_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi}_{0})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma\right\}.$ (124)
Then, the following proposition holds true.
Proposition VI.2.
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ and $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$. Then, the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ for the model (122) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in (123). Moreover this latter FIM is equal to the FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ for the model (124).
Proof:
As a direct consequence of this specific parameterisation, we immediately have that $\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})]=\mathbf{0}$ and $\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]=\mathbf{0}$. Furthermore, since $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, it follows from (119) that the FIM $\mathbf{I}_{\bm{\theta}_{0}}$ is block-diagonal, i.e. $\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}=\mathbf{0}$. Consequently the condition (121) of Proposition VI.1 is satisfied and $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ reduces to $\mathbf{I}_{\bm{\gamma}_{0}}$. ∎
VI-D The elliptical parameterized “low-rank” model
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\mathbf{0},\bm{\Sigma}_{0},g_{0})$ be a zero-mean RES-distributed vector whose scatter matrix $\bm{\Sigma}_{0}$ is modeled as:
$\bm{\Sigma}_{0}\equiv\bm{\Sigma}(\bm{\gamma}_{0},\bm{\xi}_{0})=\mathbf{A}_{0}\bm{\Xi}_{0}\mathbf{A}_{0}^{T}+\lambda_{0}\mathbf{I}_{m},$ (125)
where:
  • $\mathbf{A}_{0}\in\mathbb{R}^{m\times p}$ is a full rank column matrix with $m>p$, where $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ collects the parameter of interest and characterizes $\mathbf{A}_{0}$ and where the function $\mathbf{A}(\bm{\gamma})$ is differentiable,
  • $\bm{\Xi}_{0}\in\mathcal{S}_{p}^{\mathbb{R}}$ is a symmetric and positive definite matrix,
  • $\lambda_{0}\in\mathbb{R}^{+}$.
Consequently, the finite-dimensional nuisance vector can be defined as:
$\bm{\xi}_{0}\triangleq(\mathrm{vecs}(\bm{\Xi}_{0})^{T},\lambda_{0})^{T}\in\Psi\subset\mathrm{vecs}(\mathcal{S}_{p}^{\mathbb{R}})\times\mathbb{R}^{+}\subseteq\mathbb{R}^{r},$ (126)
with $r=p(p+1)/2+1$, while the infinite dimensional nuisance is the density generator $g_{0}\in\overline{\mathcal{G}}$.
Let us consider the two following models related to the low-rank parameterization of the scatter matrix in (125):
$\mathcal{P}_{\bm{\gamma},\bm{\xi},g}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g)=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}g(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,g\in\overline{\mathcal{G}}\right\},$ (127)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g_{0})=|\bm{\Sigma}(\bm{\gamma},\bm{\xi})|^{-1/2}g_{0}(\mathbf{x}^{T}\bm{\Sigma}(\bm{\gamma},\bm{\xi})^{-1}\mathbf{x});\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\}.$ (128)
Then, the following result can be proved by a direct application of Condition (121) of Proposition VI.1 as shown in Appendix .10.
Proposition VI.3.
For the low-rank scatter model (125), the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$.
VII Applications to circular and noncircular CES distributions
So far in this paper, we have only dealt with cases of real observation vectors. However complex-valued observations, i.e. $\mathbf{x}\in\mathbb{C}^{m}$ are an integral part of many science and engineering problems, including those in communications, radar, biomedicine, geophysics, oceanography, electromagnetics, and optics, among others. The complex field does not only provide a convenient representation for the observations but also provides a natural way to capture their physical nature as well as the transformations they go through (see e.g. [30, 31]). In many studies it has often been (implicitly) assumed that complex random vectors are circular, i.e. with invariant distribution under rotation around a center $\bm{\mu}$, that is $(\mathbf{x}-\bm{\mu})=_{d}e^{j\theta}(\mathbf{x}-\bm{\mu}),\;\forall\theta\in\mathbb{R}$ (see e.g. in [6]). This assumption however discards the information conveyed by the relationship between real and imaginary parts of the observation vectors. Consequently, the noncircularity may be an important feature that characterizes observation in many practical scenarios. For this reason, in this section, we believe it is important to provide some evidence to ensure that all results obtained so far in the case of real observations remain entirely valid for Complex Elliptically Symmetric (CES)-distributed observations. Moreover, given their importance in well-known engineering applications, we devote particular attention to the “complex” version of the Propositions VI.2 and VI.3. Before moving on, it is important to note that, although there are many different notations in the literature, in the following sections we will use the notation introduced in [8].
VII-A Real-complex representations
A random observation vector $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{C}^{m}$ is said to be CES distributed if the associated real-valued vector $\overline{\bf x}\in\mathbb{R}^{2m}$ with $\overline{\bf x}\triangleq({\rm Re}({\bf x}^{T}),{\rm Im}({\bf x}^{T}))^{T}$ is RES distributed, i.e., with pdf given by (29) when it exists. It follows that all the properties of RES distributions and propositions given for even $m$ in the previous sections apply for CES distributions. However, it is more convenient to express these properties and propositions using notations suited to the complex representation that naturally arises when using the one-to-one mapping
$\overline{\bf x}\mapsto\widetilde{\bf x}\triangleq({\bf x}^{T},{\bf x}^{H})^{T}=\sqrt{2}{\bf M}\overline{\bf x}$
where ${\bf M}\triangleq\tiny{\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}{\bf I}&i{\bf I}\\ {\bf I}&-i{\bf I}\\ \end{array}\right)}$ is a unitary matrix [8]. So, by indicating as $g_{r}\in\overline{\mathcal{G}}$ (defined in (34)) the density generator of the $2m$-dimensional real vector $\overline{\bf x}$ and if its scatter matrix $\overline{\bm{\Sigma}}\in\mathcal{S}_{2m}^{\mathbb{R}}$ is positive definite, then the pdf (29) of $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{C}^{m}$ is generally rewritten in the following form:
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},{\bf\Omega},g_{c})=|\widetilde{\bf\Sigma}|^{-1/2}g_{c}\left[2^{-1}(\widetilde{\bf x}-\widetilde{\bm{\mu}})^{H}\widetilde{\bf\Sigma}^{-1}(\widetilde{\bf x}-\widetilde{\bm{\mu}})\right],$ (129)
where
$\widetilde{\bm{\mu}}\triangleq({\bm{\mu}}^{T},{\bm{\mu}}^{H})^{T}=\sqrt{2}{\bf M}\overline{\bm{\mu}}$ (130)
with $\overline{\bm{\mu}}\in\mathbb{R}^{2m}$ denotes the location vector of $\overline{\bf x}$,
$\widetilde{\bf\Sigma}\triangleq\left(\begin{array}[]{cc}{\bf\Sigma}&{\bf\Omega}\\ {\bf\Omega}^{*}&{\bf\Sigma}^{*}\\ \end{array}\right)=2{\bf M}\overline{\bm{\Sigma}}{\bf M}^{H},\;\text{and}$ (131)
$g_{c}\triangleq 2^{m}g_{r}(2t).$ (132)
We note that ${\bf\Sigma}=E\{({\bf x}-{\bm{\mu}})({\bf x}-{\bm{\mu}})^{H}\}\in{\cal M}_{m}^{\mathbb{C}}$ and ${\bf\Omega}=E\{({\bf x}-{\bm{\mu}})({\bf x}-{\bm{\mu}})^{T}\}\in{\cal S}_{m}^{\mathbb{C}}$ where ${\cal M}_{m}^{\mathbb{C}}$ and ${\cal S}_{m}^{\mathbb{C}}$ denote the sets of all Hermitian positive definite and complex symmetric matrices, respectively.
Depending on whether ${\bf\Omega}$ is a zero-matrix or not, the CES distribution is called circular (C-CES) or non-circular (NC-CES). Due to its widespread usage, let us have a closer look at the C-CES distributions. Such particular case is characterized by structured scatter matrices $\overline{\bm{\Sigma}}=\tiny{\left(\begin{array}[]{cc}\bm{\Sigma}_{1}&-\bm{\Sigma}_{2}\\ \bm{\Sigma}_{2}&\bm{\Sigma}_{1}\end{array}\right)}$ where $\bm{\Sigma}_{1}$ and $\bm{\Sigma}_{2}$ are symmetric and skew-symmetric, respectively. Moreover, since by definition of circularity, $\bm{\Omega}=\mathbf{0}$, the C-CES pdf is a particular case of (129) that can be explicitly expressed as:
$p_{X}(\mathbf{x}|\bm{\mu},\bm{\Sigma},g_{c})=|{\bf\Sigma}|^{-1}g_{c}\left[({\bf x}-{\bm{\mu}})^{H}{\bf\Sigma}^{-1}({\bf x}-{\bm{\mu}})\right].$ (133)
Using this complex representation of even-dimensional RES distributions, the Stochastic Representation Theorem (31) can be extended to both circular and non circular CES distributions as discussed in details in [8, Sect. 1.3.2]. These “complex” stochastic representations in fact use the mutually independent, random variable $\mathcal{Q}_{c}\triangleq\frac{1}{2}\mathcal{Q}_{r}$ (where $\mathcal{Q}_{r}$ is the 2nd-order modular variate associated with $\overline{\bf x}$) and the random vector ${\bf u}_{c}\sim\mathcal{U}(S_{\mathbb{C}}^{m-1})$. Moreover, the definition of the shape matrix (37) can be straightforwardly extended while keeping definition of matrix scale function in (36). In particular for NC-CES distributions, the shape matrix
$\widetilde{\bf V}_{S}\triangleq\widetilde{\bf\Sigma}/S(\widetilde{\bf\Sigma})=\left(\begin{array}[]{cc}{\bf\Sigma}_{S}&{\bf\Omega}_{S}\\ {\bf\Omega}_{S}^{*}&{\bf\Sigma}_{S}^{*}\\ \end{array}\right)$
is structured like $\widetilde{\bf\Sigma}$.
Although certainly possible, directly rewriting all the results obtained in the previous sections using complex formalism is laborious and adds nothing to the statistical significance of our findings. Below, we therefore limit ourselves to providing two different, yet equivalent recipes that any practitioner can follow to obtain the desired complex-version of FIM and related information bounds:
Recipe 1
: Real to complex mapping.
  1. 1.
    Take the results obtained in the previous sections and consider the specific case of a real observed vector of dimension $2m$, i.e. $\mathbb{R}^{2m}\ni\overline{\bf x}\sim RES_{2m}(\overline{\bm{\mu}},\overline{\bm{\Sigma}},g_{r})$,
  2. 2.
    Use the transformations given in (130), (131) and (132) to map the “real-based” results to the “complex-based” results.
Recipe 2
: Wirtinger calculus.
  1. 1.
    Consider directly a complex observation vector, i.e. $\mathbb{C}^{m}\ni{\bf x}\sim\text{{NC-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},{\bm{\Omega}},g_{c})$ in the case of NC-CES distributions or $\mathbb{C}^{m}\ni{\bf x}\sim\text{{C-CES}}_{m}({\bm{\mu}},{\bm{\Sigma}},g_{c})$ in the case of C-CES distributions.
  2. 2.
    Use the complex Hilbert space $(\mathcal{H},\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}})$ in (1), the “complex-aware” inner product $\left\langle h_{1},h_{2}\right\rangle_{\mathcal{H}}\triangleq{E_{0}\{h_{1}h_{2}^{*}\}}$ to express projection operators and tangent spaces,
  3. 3.
    Use the Wirtinger calculus [32, 33, 34] to handle derivatives with respect to (real or/and complex) parameters.
Exstensive discussions and related examples about the above mentioned recipes can be found in [35, 36, 37, 38, 39, 24].
To conclude this section, let us now take a closer look at the “complex version” of Propositions VI.2 and VI.3 in which the parameter vector $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}$ is still assumed to be real-valued.
VII-B Applications of Proposition VI.2
Proposition VI.2 extends directly to CES distributions by considering the one-to-one mapping $\overline{\bm{\mu}}\mapsto\widetilde{\bm{\mu}}=\sqrt{2}{\bf M}\overline{\bm{\mu}}$. In fact, as shown in [40], we have:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=\beta(g_{c,0})\mathbf{J}[\widetilde{\bm{\mu}}_{0}]^{H}\widetilde{\bm{\Sigma}}_{0}^{-1}\mathbf{J}[\widetilde{\bm{\mu}}_{0}],$ (134)
for NC-CES distributions where $\beta(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m}$ with $\varphi_{c,0}(t)\triangleq\frac{-1}{g_{c,0}(t)}\frac{dg_{c,0}(t)}{dt}$ and $[\mathbf{J}[\widetilde{\bm{\mu}}_{0}]]_{i,j}\triangleq\left.\frac{\partial[\widetilde{\bm{\mu}}(\bm{\theta})]_{i}}{\partial\theta_{j}}\right|_{\bm{\theta}=\bm{\theta}_{0}}$. In the particular case of C-CES distributions, it is immediate to verify that the FIM in (134) reduces to
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\mathbf{I}_{\bm{\gamma}_{0}}=2\beta(g_{c,0}){\rm Re}\left\{\mathbf{J}[\bm{\mu}_{0}]^{H}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\right\}$ (135)
with $\mathbf{J}[\bm{\mu}_{0}]$ given by (120).
In signal processing, there are many examples of complex-valued observations ${\bf x}\in\mathbb{C}^{m}$ where the location vector $\bm{\mu}$ includes the parameters of interest $\bm{\gamma}$, while the scatter matrix $\bm{\Sigma}$ gathers the nuisance parameters $\bm{\xi}$. This is the case where a deterministic signal of interest parameterized by $\bm{\gamma}$ is disturbed by a zero-mean C-CES distributed noise, whose density generator $g$ is unspecified. A classic exemple is given by the statistical model for time delay and Doppler estimation problems (see e.g. [41, 42]).
VII-C Applications of Proposition VI.3
Proposition VI.3 also extends to C and NC-CES distributions by considering the one-to-one mapping $\overline{\bm{\Sigma}}\mapsto\widetilde{\bf\Sigma}=2{\bf M}\overline{\bm{\Sigma}}{\bf M}^{H}$. More specifically for the C-CES distribution, (125) can be rewritten in the following form [43]:
$\bm{\Sigma}_{0}=\mathbf{A}_{0}\bm{\Xi}_{0}\mathbf{A}_{0}^{H}+\lambda_{0}\mathbf{I}_{m},$ (136)
where $\bm{\Sigma}_{0}\in{\cal S}_{m}^{\mathbb{C}}$, $\bm{\Xi}_{0}\in{\cal S}_{p}^{\mathbb{C}}$ are two symmetric and positive definite matrices and $\mathbf{A}_{0}\triangleq\mathbf{A}(\bm{\gamma}_{0})\in\mathbb{C}^{m\times p}$ is full column rank that collects the parameters of interest $\bm{\gamma}_{0}$ that characterize $\mathbf{A}_{0}$.
For the C-CES distributions, Proposition VI.3 then provides the following equality where the closed-form expression of the parametric efficient FIM has been given in [40]:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(g_{c,0})}{\lambda_{0}}{\rm Re}\left\{{\bf J}[\mathrm{vec}({\bf A}_{0})]^{H}({\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{{\bf A}_{0}}){\bf J}[\mathrm{vec}({\bf A}_{0})]\right\},$ (137)
where:
${\bf H}_{0}\triangleq\bm{\Xi}_{0}{\bf A}^{H}_{0}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}\bm{\Xi}_{0},$ (138)
$\bm{\Pi}^{\bot}_{{\bf A}_{0}}\triangleq{\bf I}_{m}-{\bf A}_{0}[{\bf A}^{H}_{0}{\bf A}_{0}]^{-1}{\bf A}^{H}_{0},$ (139)
$[{\bf J}[\mathrm{vec}({\bf A}_{0})]]_{i,j}\triangleq\left.\frac{\partial[\mathrm{vec}({\bf A}(\bm{\gamma}))]_{i}}{\partial\gamma_{j}}\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ and where here $\alpha(g_{c,0})\triangleq\frac{E\{\mathcal{Q}_{c}^{2}\varphi^{2}_{c,0}(\mathcal{Q}_{c})\}}{m(m+1)}$. Note that the parametric efficient FIM (137) reduces to:
$\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}_{\textrm{DOA}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{2\alpha(g_{c,0})}{\lambda_{0}}{\rm Re}\left\{({\bf D}_{0}^{H}{\bf\Pi}^{\bot}_{{\bf A}_{0}}{\bf D}_{0})\odot{\bf H}_{0}^{T}\right\}$ (140)
for direction of arrival (DOA) modeling with one parameter per source where ${\bf A}\triangleq[{\bf a}_{1},...,{\bf a}_{p}]$ and $({\bf a}_{k})_{k=1,...,p}$ are the steering vectors parameterized by the DOA $\gamma_{k}$ with $\bm{\gamma}\triangleq(\gamma_{1},...,\gamma_{p})^{T}$ and ${\bf D}_{0}\triangleq\left.[\frac{d{\bf a}_{1}}{d\gamma_{1}},...,\frac{d{\bf a}_{p}}{d\gamma_{p}}]\right|_{\bm{\gamma}=\bm{\gamma}_{0}}$ for $p$ sources.
For the NC-CES distributions, $\widetilde{\bm{\Sigma}}_{0}$ becomes:
$\widetilde{\bm{\Sigma}}_{0}=\left(\begin{array}[]{cc}\mathbf{A}_{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{0}^{*}\end{array}\right)\widetilde{\bm{\Xi}}_{0}\left(\begin{array}[]{cc}\mathbf{A}^{H}_{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{A}_{0}^{T}\end{array}\right)+\lambda_{0}\mathbf{I}_{2m},$ (141)
where $\widetilde{\bm{\Xi}}_{0}=\left(\begin{array}[]{cc}\mathbf{\Sigma}_{c,0}&\mathbf{\Omega}_{c,0}\\ \mathbf{\Omega}_{c,0}^{*}&\mathbf{\Sigma}_{c,0}^{*}\end{array}\right)$ is structured like $\widetilde{\bm{\Sigma}}_{0}$ and (137) is also valid, where now ${\bf H}_{0}$ is replaced by [40]:
${\bf H}_{0}\triangleq\left(\mathbf{\Sigma}_{c,0}\mathbf{A}_{0}^{H},\ \mathbf{\Omega}_{c,0}\mathbf{A}_{0}^{T}\right)\widetilde{\bf\Sigma}_{0}^{-1}\left(\begin{array}[]{c}\mathbf{A}_{0}\mathbf{\Sigma}_{c,0}\\ \mathbf{A}_{0}^{*}\mathbf{\Omega}_{c,0}^{*}\end{array}\right).$ (142)
Note that (141) reduces in the so-called rectilinear case to [43]
$\widetilde{\bm{\Sigma}}_{0}=\widetilde{\mathbf{A}}_{r,0}\bm{\Xi}_{r,0}\widetilde{\mathbf{A}}_{r,0}^{H}+\lambda_{0}\mathbf{I}_{2m},$ (143)
where $\bm{\Xi}_{r,0}\in{\cal S}_{p}^{\mathbb{R}}$ is symmetric positive definite and $\widetilde{\mathbf{A}}_{r,0}=\left(\begin{array}[]{c}\mathbf{A}_{r,0}\\ \mathbf{A}_{r,0}^{*}\end{array}\right)\in\mathbb{C}^{2m\times p}$ is full column rank that collects the parameters of interest $\bm{\gamma}_{0}$ that characterize $\widetilde{\mathbf{A}}_{r,0}$. Under the condition $2m>p$, (137) is written in the form [44]:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{c,0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\frac{\alpha(g_{c,0})}{\lambda_{0}}{\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})]^{H}(\widetilde{\bf H}_{0}^{T}\otimes\bm{\Pi}^{\bot}_{\widetilde{\mathbf{A}}_{r,0}}){\bf J}[\mathrm{vec}(\widetilde{\mathbf{A}}_{r,0})],$ (144)
where $\widetilde{\bf H}_{0}\triangleq\bm{\Xi}_{r,0}\widetilde{\mathbf{A}}_{r,0}^{H}\widetilde{\bf\Sigma}_{0}^{-1}\widetilde{\mathbf{A}}_{r,0}\bm{\Xi}_{r,0}$ and $\bm{\Pi}^{\bot}_{\widetilde{\mathbf{A}}_{r,0}}\triangleq{\bf I}_{2m}-\widetilde{\mathbf{A}}_{r,0}[\widetilde{\mathbf{A}}_{r,0}^{H}\widetilde{\mathbf{A}}_{r,0}]^{-1}\widetilde{\mathbf{A}}_{r,0}^{H}$.
This low-rank scatter model encompasses many far or near-field, narrow or wide-band DOA models with scalar or vector-sensors for an arbitrary number of parameters per source and many other models such as the bandlimited SISO, SIMO [45] and MIMO [46] channel models. For example, parametrization (143) can be applied for DOA estimation modeling with rectilinear or strictly second-order sources and for SIMO channels estimation modeling with BPSK or MSK symbols [47] where $\bm{\gamma}_{0}$ represents both the localization parameters (azimuth, elevation, range) and the phase of the sources, and the real and imaginary parts of channel impulse response coefficients, respectively. Parametrization (141) on the other hand is used for DOA modeling with generally non-circular and non-rectilinear complex sources.
Remark: In all the above-mentioned applications, the observation vector ${\bf x}$ is generally assumed to be the sum of a low rank zero-mean signal of interest and a zero-mean noise term, which is mutually uncorrelated with the signal. This approach requires that both the statistical models of the useful signal and the noise are chosen a-priori. It should be noted, however, that neither the useful signal nor the noise are observable, so the chosen model could be completely misspecified. Furthermore, since they are not observable, their model cannot be estimated (in a non-parametric way) from the observed vector $\mathbf{x}$. Unlike what has been done generally in the literature, we adopt here a semi-parametric statistical model on the observed vector $\mathbf{x}$ only, without relying on any additional assumption on the statistical model on the unobservable signal of interest and noise. From a statistical point of view, we therefore believe that our approach is more valid and robust to the misspecification of the standard model generally assumed in the literature.
VIII Conclusion
The semiparametric statistical efficiency in estimation problems for elliptically symmetric distributed data was analyzed in this paper. In particular, we studied the impact of finite and infinite-dimensional nuisance parameters can have on the estimation of the parameters of interest which, in the case of elliptical distributions, are the location vector $\bm{\mu}$ and the covariance matrix $\bm{\Sigma}$. The profound and counter-intuitive result that emerged is that, in the presence of specific finite-dimensional nuisance parameters, semiparametric efficiency can be equivalent to parametric efficiency. Specifically, in the case of elliptical distributions, not knowing the density generator does not cause any loss of efficiency when estimating $\bm{\mu}$ or a scaled version of $\bm{\Sigma}$. This result had already been demonstrated by Hallin and Paindaveine using Le Cam’s asymptotic theory in [9, 10]. Unlike these works, in this article we used a purely geometric approach based on Hilbert spaces. Furthermore, as an advancement over the state of the art, we analyzed the case, important in many applications, in which the parameters of interest and the finite-dimensional nuisance parameters are given by a parameterization of the location vector and of the covariance matrix. A general condition that the parameterization in question must satisfy in order for the semiparametric efficiency to be equal to the parametric efficiency has been derived in this work. This condition therefore allows us to test this property for any particular parameterizations. Two examples were investigated here, including the well-known low-rank parameterization, often arising in many practical signal processing applications. The paper concluded with a section in which the results derived for RES distributions are extended to the case of C-CES and NC-CES distributions. The natural follow-up to this paper will be on the development of semiparametric estimators capable of achieving parametric efficiency. A promising approach for achieving this goal is that of rank-based ($R$-) estimators [48, 49, 42].
Some technical results and their proofs
  1. .1
    Some useful results in $\mathcal{H}_{q}$
    Lemma .1.
    Let $\mathcal{H}$ be an Hilbert space and let $\mathcal{U}\subseteq\mathcal{H}$ be a closed linear subspace of $\mathcal{H}$. Let us now introduce the $q$-replicating versions of $\mathcal{H}$ and $\mathcal{U}$ as $\mathcal{H}_{q}=\mathcal{H}\times\cdots\times\mathcal{H}$ and $\mathcal{H}_{q}\supseteq\mathcal{U}_{q}=\mathcal{U}\times\cdots\times\mathcal{U}$. For $\mathbf{h}\in\mathcal{H}_{q}$ and $h_{i}\in\mathcal{H}$, $i=1,\ldots,q$, we have that:
    $\left[\Pi(\mathbf{h}|\mathcal{U}_{q})\right]_{i}=\Pi(h_{i}|\mathcal{U}),\;i=1,\ldots,q,$ (145)
    where $\Pi(\mathbf{h}|\mathcal{U}_{q})\in\mathcal{U}_{q}$ and $\Pi(h_{i}|\mathcal{U})\in\mathcal{U},\;i=1,\ldots,q$.
    Proof:
    For a given $\mathbf{h}\in\mathcal{H}_{q}$, the projection $\Pi(\mathbf{h}|\mathcal{H}_{q})$ is defined as the vector in $\mathcal{H}_{q}$ such that [16, Theo. 5.1]:
    $||\mathbf{h}-\Pi(\mathbf{h}|\mathcal{H}_{q})||_{\mathcal{H}_{q}}\leq||\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}},\;\forall\mathbf{u}\in\mathcal{U}_{q}.$ (146)
    From the definition of the inner product in $\mathcal{H}_{q}$, induced by the one in $\mathcal{H}$ as $\left\langle\mathbf{h},\mathbf{u}\right\rangle_{\mathcal{H}_{q}}\triangleq\sum_{i=1}^{q}\left\langle h_{i},u_{i}\right\rangle_{\mathcal{H}}$, we have that:
    $\begin{split}||\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}}&=\sqrt{\left\langle\mathbf{h}-\mathbf{u},\mathbf{h}-\mathbf{u}\right\rangle}_{\mathcal{H}_{q}}\\ &=\sqrt{\sum\nolimits_{i=1}^{q}\left\langle h_{i}-u_{i},h_{i}-u_{i}\right\rangle_{\mathcal{H}}}=\sqrt{\sum\nolimits_{i=1}^{q}||h_{i}-u_{i}||_{\mathcal{H}}}.\end{split}$ (147)
    As a consequence, minimizing $||\mathbf{h}-\mathbf{u}||_{\mathcal{H}_{q}}$ is equivalent to minimize each term $||h_{i}-u_{i}||_{\mathcal{H}}$. Then, the equality (145) follows from the definition of orthogonal projection in $\mathcal{H}$ onto $\mathcal{U}$. ∎
    Theorem .1.
    The $q$-replicating Pythagorean theorem [13, Theo. 3.3]: Let $\mathcal{H}_{q}$ and $\mathcal{U}_{q}\subseteq\mathcal{H}_{q}$ the $q$-replicating Hilbert space and subspace defined as in Lemma .1. Let us take $\mathbf{h}\in\mathcal{H}_{q}$ and $\mathbf{u}\in\mathcal{U}_{q}$. If $\mathbf{h}\perp\mathbf{u}$, then:
    $\mathbf{G}(\mathbf{h}+\mathbf{u},\mathbf{h}+\mathbf{u})=\mathbf{G}(\mathbf{h},\mathbf{h})+\mathbf{G}(\mathbf{u},\mathbf{u}),$ (148)
    where $[\mathbf{G}(\mathbf{a},\mathbf{b})]_{i,j}\triangleq\left\langle a_{i},b_{j}\right\rangle_{\mathcal{H}}$.
    Proof:
    Let us start by showing that:
    $\mathcal{H}_{q}\ni\mathbf{h}\perp\mathbf{u}\in\mathcal{U}_{q}\Leftrightarrow\mathcal{H}\ni h_{i}\perp u_{j}\in\mathcal{U},\;i,j=1,\ldots,q.$ (149)
    1. (a)
      Proof of the implication $\Leftarrow$. Since $\mathcal{H}\ni h_{i}\perp u_{j}\in\mathcal{U}$, $i,j=1,\ldots,q$, we have that $\left\langle h_{i},u_{j}\right\rangle_{\mathcal{H}}=0$, $i,j=1,\ldots,q$. Then, $\left\langle\mathbf{h},\mathbf{u}\right\rangle_{\mathcal{H}_{q}}=\sum\nolimits_{i=1}^{q}\left\langle h_{i},u_{i}\right\rangle_{\mathcal{H}}=0$. Consequently, $\mathbf{h}\perp\mathbf{u}$.
    2. (b)
      Proof of the implication $\Rightarrow$. We notice that $\mathbf{h}\perp\mathbf{u}\Leftrightarrow\Pi(\mathbf{h}|\mathcal{U}_{q})=\mathbf{0}$. Then, from Lemma .1, $\mathbf{h}\perp\mathbf{u}\Leftrightarrow\Pi(h_{i}|\mathcal{U})=0$, $i=1,\ldots,q$ and consequently $h_{i}\perp u_{j}\in\mathcal{U}$, $i,j=1,\ldots,q$.
    Now that we know that (149) holds true, we can deduce that:
    $\left\langle\mathbf{h},\mathbf{u}\right\rangle_{\mathcal{H}_{q}}=0\Leftrightarrow\left\langle h_{i},u_{j}\right\rangle_{\mathcal{H}}=0,\;i,j=1,\ldots,q,$ (150)
    or equivalently that $\left\langle\mathbf{h},\mathbf{u}\right\rangle_{\mathcal{H}_{q}}=0\Leftrightarrow\mathbf{G}(\mathbf{h},\mathbf{u})=\mathbf{G}(\mathbf{u},\mathbf{h})=\mathbf{0}$. Finally, (148) follows from a simple calculation:
    $\begin{split}\mathbf{G}&(\mathbf{h}+\mathbf{u},\mathbf{h}+\mathbf{u})\\ &=\mathbf{G}(\mathbf{h},\mathbf{h})+\mathbf{G}(\mathbf{h},\mathbf{u})+\mathbf{G}(\mathbf{u},\mathbf{h})+\mathbf{G}(\mathbf{u},\mathbf{u})\\ &=\mathbf{G}(\mathbf{h},\mathbf{h})+\mathbf{G}(\mathbf{u},\mathbf{u}).\end{split}$ (151)
  2. .2
    Proof of Lemma II.1. From (21), it follows that $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\bar{\mathbf{s}}_{\bm{\gamma}_{0}}+\mathbf{p}$ and therefore the covariance matrix of $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ breaks down:
    $E_{0}\left\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}^{T}\right\}=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})+{E_{0}\{\mathbf{p}\mathbf{p}^{T}\}}+{E_{0}\{\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\mathbf{p}^{T}\}}+\left[{E_{0}\{\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\mathbf{p}^{T}\}}\right]^{T}.$ (152)
    Then the equality in (24) follows immediately iff $E_{0}\left\{\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\mathbf{p}^{T}\right\}=\mathbf{0}$. From the component-wise application of the inner product, we have that:
    $E_{0}\left\{\bar{\mathbf{s}}_{\bm{\gamma}_{0}}\mathbf{p}^{T}\right\}=\mathbf{0}\Leftrightarrow\left\langle[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i},p_{j}\right\rangle_{\mathcal{H}}=0\Leftrightarrow[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\perp p_{j},\ \forall i,j\in\{1,\ldots,q\}.$ (153)
    To show that $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is orthogonal to $p_{j}$, $\forall i,j$, we note that, according to its definition given in (18), $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}$ is the residual of $[\mathbf{s}_{\bm{\gamma}_{0}}]_{i}$ after projection onto $\mathcal{T}_{2}+\mathcal{T}_{3}$. As a direct consequence, we have that $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\perp(\mathcal{T}_{2}+\mathcal{T}_{3})$, or equivalently, $[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i}\in(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$. Moreover, again by definition in (25), $p_{j}\in(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}\subseteq\mathcal{T}_{2}+\mathcal{T}_{3}$. Consequently, $\left\langle[\bar{\mathbf{s}}_{\bm{\gamma}_{0}}]_{i},p_{j}\right\rangle=0$, $\forall i,j$ since $(\mathcal{T}_{2}+\mathcal{T}_{3})\perp(\mathcal{T}_{2}+\mathcal{T}_{3})^{\perp}$.
  3. .3
    Proof of Lemma II.2. Let us start by noticing that the closed subspaces $(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}$ and $\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ are orthogonal. In fact, it is immediate to verify that, each element $h\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ can be written as $h=h_{2}+h_{3}$ with $h_{2}\in\mathcal{T}_{2}$ and $h_{3}\in\mathcal{T}_{3}$. Moreover, $\forall g\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}$ we have that $g\in\mathcal{T}_{2}^{\perp}$ and $g\in\mathcal{T}_{3}^{\perp}$ and consequently $h\perp g$. In addition, $\mathcal{T}_{2}^{\perp}$ can be expressed as the (direct) sum of the and these two closed subspaces, i.e. $\mathcal{T}_{2}^{\perp}=(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp}+(\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})$. Consequently, from the property (20), we get:
    $\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|(\mathcal{T}_{2}+\mathcal{T}_{3})\cap\mathcal{T}_{2}^{\perp})=\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp})-\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}).$ (154)
    By its definition, given in (13), $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$. As a consequence, we have that the first projection in the RHS of (154) can be evaluated as $\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp})=\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$. And therefore $\mathbf{p}=\bm{0}\Leftrightarrow\bar{\mathbf{t}}_{\bm{\gamma}_{0}}=\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp})\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}\cap\mathcal{T}_{3}^{\perp}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$ and $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$. Then because $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{2}^{\perp}$, $\mathbf{p}=\bm{0}\Leftrightarrow[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\in\mathcal{T}_{3}^{\perp}$ or equivalently, $[\bar{\mathbf{t}}_{\bm{\gamma}_{0}}]_{i}\perp\mathcal{T}_{3}$.
  4. .4
    Implicit derivation of $\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$. We follow the procedure discussed in [10, Sect. 4]. Let us start by defining the mapping $v_{11}^{S}:\mathbb{R}^{m(m+1)/2-1}\mapsto\ \mathbb{R}$ implicitly defined by the constraint $S(\mathbf{V}_{S})\equiv S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$. Thanks to the implicit function theorem, under Assumptions A1, A2, A3, this mapping exists, is unique and continuously differentiable around a given $\mathrm{vecs}(\mathbf{V}_{S})$. Then, we can differentiate both side of $S(v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S})),\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=1$ w.r.t. $\underline{\mathrm{vecs}}(\mathbf{V}_{S})$ to get:
    $\begin{split}\bm{\nabla}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}S(v_{11}^{S},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))+\frac{\partial S(v_{11}^{S},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))}{\partial v_{11}^{S}}\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=\mathbf{0},\end{split}$ (155)
    then consequently
    $\begin{split}\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}v_{11}^{S}(\underline{\mathrm{vecs}}(\mathbf{V}_{S}))=-\frac{\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}S(v_{11}^{S},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))}{\partial S(v_{11}^{S},\underline{\mathrm{vecs}}(\mathbf{V}_{S}))/\partial v_{11}^{S}}.\end{split}$ (156)
  5. .5
    Explicit calculation of $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}$ for $S(\bm{\Sigma})=[\bm{\Sigma}]_{1,1}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$. For the scale $S(\bm{\Sigma})=[\bm{\Sigma}]_{1,1}$ it is trivial to verify that $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=\mathbf{0}^{T}_{m(m+1)/2-1}$. For the scale $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$, let us start by noticing that it implies $\mathrm{tr}(\mathbf{V}_{S})=m$. By taking the differential on both side of this equality, and by expliting the linearity of the trace, we get:
    $d\mathrm{tr}(\mathbf{V}_{S})=\mathrm{tr}(\mathbf{I}_{m}d\mathbf{V}_{S})=\mathrm{vec}(\mathbf{I}_{m})^{T}\mathrm{vec}(d\mathbf{V}_{S})=0.$ (157)
    Now, by definition of $\mathbf{D}_{m}$:
    $\mathrm{vec}(\mathbf{V}_{S})=\mathbf{D}_{m}\mathrm{vec}(\mathbf{V}_{S})=\mathbf{D}_{m}\left(\begin{array}[]{c}[\mathbf{V}_{S}]_{11}\\ \underline{\mathrm{vecs}}(\mathbf{V}_{S})\end{array}\right).$ (158)
    Then, by substituting (158) in (157), we obtain:
    $\begin{split}\mathrm{vec}(\mathbf{I}_{m})^{T}&\mathbf{D}_{m}\left(\begin{array}[]{c}d[\mathbf{V}_{S}]_{11}\\ d\underline{\mathrm{vecs}}(\mathbf{V}_{S})\end{array}\right)=(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}d[\mathbf{V}_{S}]_{11}+\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}d\underline{\mathrm{vecs}}(\mathbf{V}_{S})=0,\end{split}$ (159)
    from which we immediately get:
    $\left(\frac{d[\mathbf{V}_{S}]_{11}}{d\underline{\mathrm{vecs}}(\mathbf{V}_{S})}\right)^{T}\triangleq\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{I}_{m})^{T}\mathbf{D}_{m})_{1}}.$ (160)
    The scale $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$ implies that $|\mathbf{V}_{S}|=1$. Taking as before the differential on both side, we have [23, pg. 149]:
    $|\mathbf{V}_{S}|\mathrm{tr}(\mathbf{V}_{S}^{-1}d\mathbf{V}_{S})=0\Rightarrow\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathrm{vec}(d\mathbf{V}_{S})=0.$ (161)
    Consequently, by applying exactly the same procedure as before, we get:
    $\nabla_{\underline{\mathrm{vecs}}(\mathbf{V}_{S})}^{T}[\mathbf{V}_{S}]_{11}=-\frac{\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m}\underline{\mathbf{I}}_{m}^{T}}{(\mathrm{vec}(\mathbf{V}_{S}^{-1})^{T}\mathbf{D}_{m})_{1}}.$ (162)
  6. .6
    Properties of the matrix $\mathbf{M}_{S}^{\bm{\Sigma}}$. Following the notation in [10], let us define the matrix $\mathbf{D}_{S}^{\bm{\Sigma}}$ as the matrix derivative of the scale function $S$ defined in (36):
    $\mathbf{D}_{S}^{\bm{\Sigma}}\triangleq\frac{\partial S(\bm{\Sigma})}{\partial\bm{\Sigma}}\in\mathbb{R}^{m\times m}.$ (163)
    We note that, for $S(\bm{\Sigma})=[\bm{\Sigma}]_{11}$, $S(\bm{\Sigma})=\mathrm{tr}(\bm{\Sigma})/m$ and $S(\bm{\Sigma})=|\bm{\Sigma}|^{1/m}$, we have $\mathbf{D}_{S}^{\bm{\Sigma}}=\mathbf{e}_{1}\mathbf{e}_{1}^{T}$, $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}\mathbf{I}_{m}$ and $\mathbf{D}_{S}^{\bm{\Sigma}}=m^{-1}|\bm{\Sigma}|^{1/m}\bm{\Sigma}^{-1}$ respectively.
    The matrices $\mathbf{D}_{S}^{\bm{\Sigma}}$ and $\mathbf{M}_{S}^{\bm{\Sigma}}$ have the following properties [10]:
    • P1
      From the 1-homogeneity of $S$, we have that $\mathbf{D}_{S}^{c\bm{\Sigma}}=\mathbf{D}_{S}^{\bm{\Sigma}}$ for all $c>0$. Moreover $S(\bm{\Sigma})=\mathrm{tr}(\mathbf{D}_{S}^{\bm{\Sigma}}\bm{\Sigma})$.
    • P2
      Let $\mathbf{A}$ be a $m\times m$ symmetric, real matrix. If $\mathrm{tr}(\mathbf{D}_{S}^{\bm{\Sigma}}\mathbf{A})=0$, then
      $[\mathbf{M}_{S}^{\bm{\Sigma}}]^{T}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathbf{K}_{\mathbf{V}_{S,0}}\underline{\mathrm{vecs}}(\mathbf{A})=\mathbf{D}_{m}\mathrm{vecs}(\mathbf{A})=\mathrm{vec}(\mathbf{A}).$ (164)
    • P3
      $\mathbf{M}_{S}^{\bm{\Sigma}}$ has full row rank equal to $m(m+1)/2-1$.
    • P4
      $\text{Ker}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\bm{\Sigma}}),a\in\mathbb{R}\}$, that is the null space of the restriction to the vectorized space $\mathcal{S}_{m}^{\mathbb{R}}$ of the real, symmetric matrices of dimension $m\times m$ of the linear application defined by $\mathbf{M}_{S}^{\bm{\Sigma}}$ is the one-dimensional space generated by $\mathrm{vec}(\mathbf{D}_{S}^{\bm{\Sigma}})$.
    The proof of these properties can be found in [10, Lemma 4.2]. Let us now focus on the property P4 to add some insight on the image of $\mathbf{M}_{S}^{\bm{\Sigma}}$. From the Property P3, we have that the dimension of $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$, i.e. the number of the linearly independent rows of $\mathbf{M}_{S}^{\bm{\Sigma}}$ is equal to $m(m+1)/2-1$. Moreover, as a direct consequence of the rank-nullity theorem, we have the following additional property:
    • P5
      An orthonormal basis of the $m(m+1)/2-1$-dimensional image $\text{Im}\;\mathbf{M}_{S}^{\bm{\Sigma}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the restriction to $\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ of the linear application defined by $\mathbf{M}_{S}^{\bm{\Sigma}}$ is given by the columns of the matrix $\mathbf{V}_{\bm{\Sigma}}=\mathbf{D}_{m}\mathbf{U}_{\bm{\Sigma}}\in\mathbb{R}^{m^{2}\times(m(m+1)/2-1)}$ such that :
      $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}})^{T}\mathbf{U}_{\bm{\Sigma}}=\mathbf{0},\quad\mathbf{U}_{\bm{\Sigma}}^{T}\mathbf{U}_{\bm{\Sigma}}=\mathbf{I}_{m(m+1)/2-1}.$ (165)
      From its definition given in (55), we have that $\left[\mathbf{M}_{S}^{\bm{\Sigma}}\right]^{T}=\mathbf{D}_{m}\mathbf{K}_{\bm{\Sigma}}$. Consequently the column of the matrix $\mathbf{K}_{\bm{\Sigma}}$ and the one of $\mathbf{U}_{\bm{\Sigma}}$ span the same subspace. In other word, for each $\bm{\Sigma}\in\mathcal{S}_{m}^{\mathbb{R}}$, there exists an invertible transformation matrix $\mathbf{S}_{\bm{\Sigma}}$ such that:
      $\mathbf{K}_{\bm{\Sigma}}\mathbf{S}_{\bm{\Sigma}}=\mathbf{U}_{\bm{\Sigma}}.$ (166)
      We note, for further reference, that for $\mathbf{V}=s^{-1}\bm{\Sigma}$ from the Property P1, we have $\mathrm{vecs}(\mathbf{D}_{S}^{\bm{\Sigma}})^{T}\mathbf{U}_{\bm{\Sigma}}=\mathrm{vecs}(\mathbf{D}_{S}^{\mathbf{V}})^{T}\mathbf{U}_{\mathbf{V}}$ and then $\mathbf{U}_{\mathbf{V}}=\mathbf{U}_{\bm{\Sigma}}$ and consequently:
      $\mathbf{K}_{\mathbf{V}}\mathbf{S}_{\mathbf{V}}=\mathbf{U}_{\mathbf{V}}.$ (167)
    Remarkably, the matrix $\mathbf{U}_{\bm{\Sigma}}$ is the same matrix defined in [22, Sect. IV.D] in a different set-up and it will be the bridge to prove some interesting equalities.
  7. .7
[…]
    $\mathcal{P}_{g}=\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g):g\in\bar{\mathcal{G}}\right\}.$
    Clearly, this model is the restriction to the RES-distribution of the general non-parametric model in (9) introduced in Sect. II. The aim of this appendix is to derive the related tangent space $\mathcal{T}_{g_{0}}$ (c.t. $\mathcal{T}_{3}$ in Sect. II). To this end, following [13, Sec. 4.2], [1, Sec. 3.1],[2, Sec. 2.2], we need to introduce the set of parametric sub-models of $\mathcal{P}_{g}$. Formally, the i-th parametric sub-model of $\mathcal{P}_{g}$ is defined as:
    $\mathcal{P}_{\tau_{\bm{\rho},i}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho},i}),\bm{\rho}\in\Upsilon_{i}\subseteq\mathbb{R}^{r_{i}}\right\},$ (168)
    where:
    $\begin{split}\tau_{\bm{\rho},i}:\,&\mathcal{X}\times\Upsilon_{i}\rightarrow\bar{\mathcal{G}}\\ &\bm{\rho}\mapsto\tau_{i}(\mathbf{x},\bm{\rho}),\end{split}$ (169)
    is a known function parametrized by an unknown finite-dimensional vector $\bm{\rho}$. In particular, for every $i\in\mathbb{N}$, $\mathcal{P}_{\bm{\theta},\tau_{i}}$ in (168) is a parametric model satisfying the following three conditions [13, Sec. 4.2]:
    C0)
    $\tau_{\bm{\rho},i}:\mathcal{X}\times\Upsilon_{i}\rightarrow\bar{\mathcal{G}}$ is a smooth parametric map $\forall i\in\mathbb{N}$,
    C1)
    $\mathcal{P}_{\tau_{i}}\subseteq\mathcal{P}_{g},\;\forall i\in\mathbb{N}$,
    C2)
    $p_{0}(\mathbf{x})\in\mathcal{P}_{\tau_{i}}$, i.e. $\forall i\in\mathbb{N}$ there exists a vector $\bm{\rho}_{0}\in\Upsilon_{i}$ such that $p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})=p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},g_{0})$.
    Intuitively, a parametric sub-model $\mathcal{P}_{\tau_{i}}$ can be though as a finite-dimensional approximation of the non-parametric model $\mathcal{P}_{g}$. The purpose of using a parametric sub-model lies in the fact that its tangent space is well-defined as shown in (12) as:
    $\mathcal{H}\supseteq\mathcal{T}_{\bm{\rho}_{0},i}\triangleq\mathrm{Span}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{1},\ldots,[\mathbf{s}_{\bm{\rho}_{0},i}]_{r_{i}}\},$ (170)
    where $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})$ is the score vector of $\bm{\rho}_{0}\in\Upsilon_{i}$ in the $i-th$ parametric sub-model $\mathcal{P}_{\tau_{i}}$. Consequently, according to [14] and [13, Sec. 4.4], the tangent space $\mathcal{T}_{g_{0}}$ can be defined as the closure  [^4] The closure $\overline{\mathcal{A}}$ of a set $\mathcal{A}$ is defined as the smallest closed set that contains $\mathcal{A}$, or equivalently, as the set of all elements in $\mathcal{A}$ together with all the limit points of $\mathcal{A}$. of the union of all the (parametric) tangent spaces $\mathcal{P}_{\tau_{i}}$:
    $\mathcal{T}_{g_{0}}=\overline{\bigcup\nolimits_{i\in\mathbb{N}}\mathcal{T}_{\bm{\rho}_{0,i}}}\subseteq\mathcal{H}.$ (171)
    Equivalently, $\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ is the subspace  [^5] The closure of a union of linear spaces doesn’t need to be linear, in general. However, as discussed in [50, Assumption S] and [13, Sec. 4.4, Remark 5]), $\mathcal{T}_{g_{0}}$ is a linear space in the vast majority of the non-pathological statistical models. of $\mathcal{H}$ composed by all the functions $h\in\mathcal{T}_{g_{0}}\subseteq\mathcal{H}$ for which there exists a sequence $\{\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}\}_{i\in\mathbb{N}}$ such that $||h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}}||^{2}=E_{0}\left\{(h-\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0,i}})^{2}\right\}\rightarrow 0$.
    Now that we have the theoretical and formal framework, let us go back to the application at hand. Specifically, we have to show that $\mathcal{T}_{g_{0}}$ can actually be expressed as in (75):
    $\begin{split}\mathcal{T}_{g_{0}}=\left\{h\in\mathcal{H}|h(\mathbf{x})=h(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})),a.s.\;\mathbf{x}\in\mathcal{X}\right\},\end{split}$ (172)
    where, from (32), $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})=s_{0}^{-1}(\mathbf{x}-\bm{\mu}_{0})^{T}\mathbf{V}_{S,0}^{-1}(\mathbf{x}-\bm{\mu}_{0})$.
    Proof:
    The proof follows from some small modifications of the one in [13, Theo. 4.4]. Specifically, we need to show that:
    • i)
      Any element of $\mathcal{T}_{\bm{\rho}_{0},i},\forall i\in\mathbb{N}$ is an element of $\mathcal{T}_{g_{0}}$, i.e. $\mathcal{T}_{\bm{\rho}_{0},i}\subset\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$,
    • ii)
      Any element of $\mathcal{T}_{g_{0}}$ can be expressed as an element of a given $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, for some $\tilde{i}\in\mathbb{N}$, or as a converging sequence of such elements, i.e. $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ where $\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ is the closure in $\mathcal{H}$ of $\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}$.
    Let us start by showing i). Each element of $\mathcal{T}_{\bm{\rho}_{0},i},\forall i\in\mathbb{N}$ is of the form $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i},\mathbf{c}_{i}\in\mathbb{R}^{r_{i}}$ where:
    $\mathbf{s}_{\bm{\rho}_{0},i}=\nabla_{\bm{\rho}}\ln\left[s_{0}^{-m/2}|\mathbf{V}_{S,0}|^{-1/2}\tau_{\bm{\rho}_{0,i}}\left(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})\right)\right].$ (173)
    Consequently, $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}$ is clearly an element of $\mathcal{T}_{g_{0}}$ thanks to the linearity of the expectation operator and to the fact that $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\}=0$ and $E_{0}\{[\mathbf{s}_{\bm{\rho}_{0},i}]^{2}_{j}\}<+\infty$, for $j=1,\ldots,r_{i}$ from usual properties of the score vectors (under the regularity conditions discussed in [17, Sects. 6.2, 6.3]). Then $[\mathbf{s}_{\bm{\rho}_{0},i}]_{j}\in\mathcal{H}$ for $j=1,\ldots,r_{i}$ and it is measurable w.r.t. $Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x})$, then $\mathbf{c}_{i}^{T}\mathbf{s}_{\bm{\rho}_{0},i}\in\mathcal{T}_{g_{0}},\forall i\in\mathbb{N}$.
    Let us now move to ii). Let star by choosing $r_{\tilde{i}}$ elements $\{\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\}_{j=1}^{r_{\tilde{i}}}\in\mathcal{T}_{g_{0}}$ such that $\tilde{h}_{j}$ are bounded functions. Then, as parametric sub-model of the form in (168), we may choose the following one:
    $\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}=\left\{p_{X}(\mathbf{x}|\bm{\rho})=p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\left[1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right]\right\}.$ (174)
    where $\bm{\rho}\in\Upsilon_{\tilde{i}}$ is sufficiently small to guarantee that
    $1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\geq 0,\forall\mathbf{x}\in\mathcal{X},$ (175)
    such then $p_{X}(\mathbf{x}|\bm{\rho})\geq 0$. Note that, such “small $\bm{\rho}$” exists since we are working with bounded functions $\tilde{h}_{j}$. Moreover, for each $p_{X}(\mathbf{x}|\bm{\rho})\in\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}$, we have that:
[…]
    since $\tilde{h}_{j}\in\mathcal{H}$, then $p_{X}(\mathbf{x}|\bm{\rho})$ is a proper pdf and consequently $\mathcal{P}_{\tau_{\bm{\rho},\tilde{i}}}$ is a proper parametric sub-model that satisfies the condition C0, C1 and C2. Now, a score vector of this specific parametric sub-model is of the form:
    $\begin{split}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}&=\nabla_{\bm{\rho}}\ln p_{X}(\mathbf{x}|\bm{\gamma}_{0},s_{0},\tau_{\bm{\rho}_{0},i})=\\ &=\left.\nabla_{\bm{\rho}}\ln\left[p_{0}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\left(1+\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right)\right]\right|_{\bm{\rho}=\bm{\rho}_{0}}\\ &=\left.\nabla_{\bm{\rho}}\sum\nolimits_{j=1}^{r_{\tilde{i}}}\rho_{j}\tilde{h}_{j}\right|_{\bm{\rho}=\bm{\rho}_{0}}=(\tilde{h}_{1},\ldots,\tilde{h}_{r_{\tilde{i}}})^{T}.\end{split}$ (177)
    Consequently, since, as said before, any element of $\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$ is of the form $\mathbf{c}^{T}\mathbf{s}_{\bm{\rho}_{0},\tilde{i}}$ for some $\mathbf{c}\in\mathbb{R}^{r_{\tilde{i}}}$, we just need to choose $\mathbf{c}=\mathbf{e}_{j}$ to prove that $\mathcal{T}_{g_{0}}\ni\tilde{h}_{j}(Q_{\bm{\mu}_{0},s_{0}\mathbf{V}_{S,0}}(\mathbf{x}))\in\mathcal{T}_{\bm{\rho}_{0},\tilde{i}}$, where $\tilde{h}_{j}$ is a bounded function. The proof is completed by noticing that the set of bounded functions is dense in $\mathcal{H}$ and consequently, any element $h\in\mathcal{T}_{g_{0}}$ can be obtained as a converging sequence of bounded functions. This allow us to state that $\mathcal{T}_{g_{0}}\subset\overline{\mathcal{T}_{\bm{\rho}_{0,\tilde{i}}}}$ and conclude the proof. ∎
  8. .8
    Explicit expressions of the CRBs in the models $\mathcal{P}_{\bm{\eta},g_{0}}$ and $\mathcal{P}_{\bm{\nu},g_{0}}$. Let us first evaluate the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma}_{0})^{T})^{T}\in\Omega$ in (41) as the inverse of the related FIM $\mathbf{I}_{\bm{\nu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\nu}_{0}}\mathbf{s}_{\bm{\nu}_{0}}^{T}\}$. To this end, we note that the two components of the related score vector $\mathbf{s}_{\bm{\nu}_{0}}=(\mathbf{s}_{\bm{\mu}_{0}}^{T},\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T})^{T}$ have been already introduced in (56) and (57). Consequently, from standard calculation and by using the independence between $\mathcal{Q}$ and $\mathbf{u}$ (along with the properties of $\mathbf{u}$), the FIM for $\bm{\nu}_{0}\in\Omega$ in $\mathcal{P}_{\bm{\nu},g_{0}}$ is given by:
    $\mathbf{I}_{\bm{\nu}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\nu}_{0}}\mathbf{s}_{\bm{\nu}_{0}}^{T}\}=\left(\begin{array}[]{cc}\mathbf{I}_{\bm{\mu}_{0}}&\mathbf{I}_{\bm{\mu}_{0},\mathrm{vecs}(\bm{\Sigma}_{0})}\\ \mathbf{I}_{\bm{\mu}_{0},\mathrm{vecs}(\bm{\Sigma}_{0})}^{T}&\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\\ \end{array}\right),$ (178)
    where
    $\mathbf{I}_{\bm{\mu}_{0}}\triangleq E_{0}\left\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\bm{\mu}_{0}}^{T}\right\}=\beta(g_{0})\bm{\Sigma}_{0}^{-1},$ (179)
    $\mathbf{I}_{\bm{\mu}_{0},\mathrm{vecs}(\bm{\Sigma}_{0})}\triangleq E_{0}\left\{\mathbf{s}_{\bm{\mu}_{0}}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T}\right\}=\mathbf{0},$ (180)
    $\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\triangleq E_{0}\left\{\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}\mathbf{s}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{T}\right\}=\mathbf{D}_{m}^{T}\left[\frac{1}{2}\alpha(g_{0})(\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1})+\frac{1}{4}({\alpha(g_{0})-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{D}_{m}.$ (181)
    Then the CRB on $\bm{\nu}_{0}$ in (41) can be derived from the block diagonal structured FIM in (178) as:
    $\mathrm{CRB}(\bm{\nu}_{0})\triangleq\mathbf{I}_{\bm{\nu}_{0}}^{-1}=\left(\begin{array}[]{cc}\mathrm{CRB}(\bm{\mu}_{0})&\mathbf{0}\\ \mathbf{0}&\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\\ \end{array}\right),$ (182)
    where:
    $\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\beta(g_{0})^{-1}\bm{\Sigma}_{0}\ \ \mbox{and}\ \ \mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\triangleq\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{-1}.$ (183)
    To calculate $\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}^{-1}$, let us rewrite (181) in the form $\mathbf{I}_{\mathrm{vecs}(\bm{\Sigma}_{0})}=\mathbf{D}_{m}^{T}[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]\mathbf{D}_{m}$. The inverse of the middle term of (181) can be derived from the inversion matrix lemma giving:
    $[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}=2\alpha^{-1}(g_{0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (184)
    Then using $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}=\frac{1}{2}(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})$ and $\mathbf{K}_{m}(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})=(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\mathbf{K}_{m}$ [23, Ch. 3], we straightforwardly get
    $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}[\mathbf{A}+\mathbf{a}\mathbf{a}^{T}]^{-1}\mathbf{D}_{m}^{\#T}\mathbf{D}^{T}_{m}=\alpha^{-1}(g_{0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{2\alpha^{-1}(g_{0})(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}.$ (185)
    Finally, using $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$, $\mathbf{D}_{m}^{\#}\mathbf{D}_{m}=\mathbf{I}_{m(m+1)/2}$, [23, Ch. 3], it is easy to check that
[…]
    and thus, since $\mathbf{D}_{m}^{\#}{\mathbf{K}}_{m}=\mathbf{D}_{m}^{\#}$ [23, Ch. 3]
    $\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))=2\alpha^{-1}(g_{0})\mathbf{D}_{m}^{\#}\left[(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})-\frac{(\alpha(g_{0})-1)}{(m+2)\alpha(g_{0})-m}\mathrm{vec}(\bm{\Sigma}_{0})\mathrm{vec}(\bm{\Sigma}_{0})^{T}\right]\mathbf{D}_{m}^{\#T}.$ (187)
    In the second step of the proof, the CRB on $\bm{\eta}_{0}$ is given by
    $\mathrm{CRB}(\bm{\eta}_{0})=\mathbf{J}[\mathbf{w}^{-1}](\bm{\nu}_{0})\mathrm{CRB}(\bm{\nu}_{0})[\mathbf{J}[\mathbf{w}^{-1}](\bm{\nu}_{0})]^{T},$ (188)
    from the inverse diffeomorphism of $\mathbf{w}$ (50), whose Jacobian matrix $\mathbf{J}[\mathbf{w}^{-1}](\bm{\nu}_{0})$ is given by (95).
    Let us explicitly evaluate the term $\nabla_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma}_{0})$ in (95). From (37), we have
    $\nabla_{\mathrm{vec}(\bm{\Sigma})}^{T}\mathrm{vec}(\mathbf{V}_{S})=S^{-1}(\bm{\Sigma})\left[\mathbf{I}_{m^{2}}-\mathrm{vec}(\mathbf{V}_{S})\nabla_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma})\right].$ (189)
    Then from $\underline{\mathrm{vecs}}(\mathbf{V}_{S})=\underline{\mathbf{I}}_{m}\mathrm{vecs}(\mathbf{V}_{S})$ and $\mathrm{vec}(\bm{\Sigma})=\mathbf{D}_{m}\mathrm{vecs}(\bm{\Sigma})$, we get:
    $\nabla_{\mathrm{vecs}(\bm{\Sigma})}^{T}\underline{\mathrm{vecs}}(\mathbf{V}_{S})=S^{-1}(\bm{\Sigma})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\left[\mathbf{I}_{m^{2}}-\mathrm{vec}(\mathbf{V}_{S})\nabla_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma})\right]\mathbf{D}_{m}$ (190)
    Consequently, the CRB on $\bm{\eta}_{0}$ is given by
    $\mathrm{CRB}(\bm{\eta}_{0})=\left(\begin{array}[]{ccc}\mathrm{CRB}(\bm{\mu}_{0})&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})&\bm{\Psi}\\ \mathbf{0}&\bm{\Psi}^{T}&\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}))\\ \end{array}\right),$ (191)
    where $\mathrm{CRB}(\bm{\mu}_{0})$ is given by (97) and:
    $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})$ $=$ $\nabla_{\mathrm{vecs}(\bm{\Sigma})}^{T}\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))[\nabla_{\mathrm{vecs}(\bm{\Sigma})}\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})],$ (192)
    $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}))$ $=$ $\nabla_{\mathrm{vecs}(\bm{\Sigma})}^{T}S(\bm{\Sigma}_{0})\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\nabla_{\mathrm{vecs}(\bm{\Sigma})}S(\bm{\Sigma}_{0}),$ (193)
    $\bm{\Psi}$ $=$ $\nabla_{\mathrm{vecs}(\bm{\Sigma})}^{T}\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})\mathrm{CRB}(\mathrm{vecs}(\bm{\Sigma}_{0}))\nabla_{\mathrm{vecs}(\bm{\Sigma})}S(\bm{\Sigma}_{0}).$ (194)
    Then, applying Euler’s homogeneous function theorem to the score function $S(\bm{\Sigma})$ which is homogeneous of order one:
    $\nabla_{\mathrm{vec}(\bm{\Sigma})}^{T}S(\bm{\Sigma})\mathrm{vec}(\bm{\Sigma})=S(\bm{\Sigma}),$ (195)
    the relation $\mathbf{D}_{m}\mathbf{D}_{m}^{\#}=\frac{1}{2}(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})$ and the fact that $\mathbf{K}_{m}\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0})=\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0})$ the relations (192), (193) and (194) can be explicitly expressed as:
    $\mathrm{CRB}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})=\alpha^{-1}(g_{0})\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S}^{T}(\mathbf{V}_{S,0})\mathbf{D}_{m}^{\#T}\underline{\mathbf{I}}_{m}^{T},$ (196)
    that correspond to the expression reported in (98) and:
    $\mathrm{CRB}(s_{0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S,0}))=\frac{2s_{0}^{2}}{\alpha(g_{0})}\left[\nabla_{{\rm vec}({\bm{\Sigma})}}^{T}S(\bm{\Sigma}_{0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0})-\left(\frac{\alpha(g_{0})-1}{(m+2)\alpha(g_{0})-m}\right)\right],$ (197)
    $\bm{\Psi}=2\alpha^{-1}(g_{0})s_{0}\underline{\mathbf{I}}_{m}\mathbf{D}_{m}^{\#}\mathbf{P}_{S}(\mathbf{V}_{S,0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S(\bm{\Sigma}_{0}),$ (198)
    Remarkably, it can be shown that, for the scale functional $S_{d}(\bm{\Sigma})\triangleq|\bm{\Sigma}|^{1/m}$ the term $\bm{\Psi}$ cancels out, in accordance with Proposition IV.1. In fact, for $S_{d}(\bm{\Sigma})$, we get $\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0}^{-1})$ from [23, Ch. 8, Th. 1] and $\mathbf{P}_{S_{d}}(\mathbf{V}_{S,0})=\mathbf{I}_{m^{2}}-\frac{1}{m}{\rm vec}(\bm{\Sigma}_{0}){\rm vec}(\bm{\Sigma}_{0}^{-1})^{T}$. Consequently $(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0})$ and thus
    $\mathbf{P}_{S_{d}}(\mathbf{V}_{S,0})(\bm{\Sigma}_{0}\otimes\bm{\Sigma}_{0})\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}}{m}{\rm vec}(\bm{\Sigma}_{0})-\frac{s_{0}}{m^{2}}\mathrm{tr}({\mathbf{I}}_{m}){\rm vec}(\bm{\Sigma}_{0})=\mathbf{0}$ and therefore $\bm{\Psi}=\mathbf{0}$.
    Moreover, $\mathbf{P}_{S_{d}}(\mathbf{V}_{S,0})(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\mathbf{P}_{S_{d}}^{T}(\mathbf{V}_{S,0})=(\mathbf{I}_{m^{2}}+\mathbf{K}_{m})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})-\frac{2}{m}{\rm vec}(\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0})^{T}$ using in particular ${\rm vec}(\mathbf{V}_{S,0}^{-1})^{T}(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0}^{-1})=\mathrm{tr}(\mathbf{V}_{S,0}^{-1}\mathbf{V}_{S,0}\mathbf{V}_{S,0}^{-1}\mathbf{V}_{S,0})=m$, which proves (99).
    Finally, $\nabla_{{\rm vec}({\bm{\Sigma})}}^{T}{S_{d}}(\bm{\Sigma}_{0})(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0})\nabla_{{\rm vec}({\bm{\Sigma})}}S_{d}(\bm{\Sigma}_{0})=\frac{s_{0}^{2}}{m^{2}}{\rm vec}(\mathbf{V}_{S,0}^{-1})^{T}(\mathbf{V}_{S,0}\otimes\mathbf{V}_{S,0}){\rm vec}(\mathbf{V}_{S,0}^{-1})=\frac{s_{0}^{2}}{m}$, which proves that (197) reduces to:
    $\mathrm{CRB}(s_{d,0}|\underline{\mathrm{vecs}}(\mathbf{V}_{S_{d},0}))=\frac{4|\bm{\Sigma}|^{2/m}}{m[m(\alpha(g_{0})-1)+2\alpha(g_{0})]}.$ (199)
  9. .9
    Proof of Proposition VI.1. Let us apply Lemma II.2 to the case of RES distributions for which $\mathcal{T}_{3}=\mathcal{T}_{g_{0}}$. From the definition of $\mathcal{T}_{g_{0}}$ in (75) with the related orthogonal projection in (76), Condition (28) of Lemma II.2 can be expressed as:
    $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})|\mathcal{Q}\right\}=\mathbf{0}.$ (200)
    Then, by using the expression of $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{2})$ deduced from (14), we have:
    $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{Q}\right\}=\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\},$ (201)
    where $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{Q}\right\}$ and $E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\}$ the two sub-vector of $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$. To conclude the proof, we just need to evaluate $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$. To this end, from the expression of $\mathbf{s}_{\bm{\theta}_{0}}$ given in (116) and by noticing that:
    $\mathrm{tr}\left(\mathbf{P}_{i}^{0}\right)=\mathrm{tr}\left(\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1}\right)=\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\mathrm{vec}(\bm{\Sigma}_{i}^{0}),$ (202)
    we immediately have that:
    $\begin{split}E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}&=_{d}-2^{-1}(1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q}))\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\\ &=-2^{-1}(1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q}))[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1}).\end{split}$ (203)
    Consequently, by substituting (203) in (201) and by noticing that $\mathrm{Pr}\left(\{1+m^{-1}\mathcal{Q}\psi_{0}(\mathcal{Q})=0\}\right)=0$ (since $\mathcal{Q}$ is a continuous random variable), we get (121).
  10. .10
    Proof of Proposition VI.3. The different steps of the proof are based on some reasoning and notations of [44, Sec. VI of supplement material] for C-CES distributions and the general ”low-rank” model (137), that itself takes up the steps of the proof presented in [51].
    We deduce from the FIM in $\mathbf{I}_{\bm{\theta}_{0}}$ (119), the following expressions of the sub-blocks:
    $\mathbf{I}_{\bm{\xi}_{0}}$ $=$ $\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2}){\bf T}^{1/2}\right]\left[{\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\right]\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]$ (204)
    $\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}$ $=$ $\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2}){\bf T}^{1/2}\right]\left[{\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\right]\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]$ (205)
    with
    ${\bf T}\triangleq 2^{-1}\alpha(g_{0}){\bf I}_{m^{2}}+4^{-1}(\alpha(g_{0})-1)\mathrm{vec}({\bf I}_{m})\mathrm{vec}({\bf I}_{m})^{T},$ (206)
    which can be written, by means of the notation used in [44, Sec; V of supplement material], in the following compact form
    $\mathbf{I}_{\bm{\xi}_{0}}={\bm{\Delta}}^{T}{\bm{\Delta}}\ \ \mbox{and}\ \ \mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}={\bf G}^{T}{\bm{\Delta}},$ (207)
    with
    ${\bm{\Delta}}\triangleq\left[{\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\right]\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]\ \ \mbox{and}\ \ {\bf G}\triangleq\left[{\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\right]\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})].$ (208)
    Consequently the left hand side of (121) can be expressed as:
    ${\bf G}^{T}\left[{\bf I}_{m^{2}}-{\bm{\Delta}}({\bm{\Delta}}^{T}{\bm{\Delta}})^{-1}{\bm{\Delta}}^{T}\right]{\bf T}^{-1/2}(\bm{\Sigma}_{0}^{1/2}\otimes\bm{\Sigma}_{0}^{1/2})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})={\bf G}^{T}{\bm{\Pi}}^{\bot}_{\bm{\Delta}}{\bf T}^{-1/2}\mathrm{vec}({\bf I}_{m}),$ (209)
    with ${\bm{\Pi}}^{\bot}_{\bm{\Delta}}\triangleq{\bf I}_{m^{2}}-{\bm{\Delta}}({\bm{\Delta}}^{T}{\bm{\Delta}})^{-1}{\bm{\Delta}}^{T}$. This implies that condition (121) of Proposition VI.1 is satisfied iif
    $\mathrm{vec}({\bf I}_{m})^{T}{\bf T}^{-1/2}{\bm{\Pi}}^{\bot}_{\bm{\Delta}}{\bf g}_{k}=0,\ \ k=1,..,q,$ (210)
    where ${\bf g}_{k},k=1,..,q$ denotes the $k$th column of ${\bf G}$. Let’s further partition the matrix ${\bm{\Delta}}$ as
    ${\bm{\Delta}}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\left[\mathbf{J}_{\mathrm{vecs}(\bm{\Xi})}[\mathrm{vec}(\bm{\Sigma}_{0})]\ |\ \mathbf{J}_{\lambda}[\mathrm{vec}(\bm{\Sigma}_{0})]\right]\triangleq[{\bf V}\ |\ {\bf u}_{n}],$ (211)
    with ${\bf u}_{n}={\bf T}^{1/2}(\bm{\Sigma}_{0}^{-1/2}\otimes\bm{\Sigma}_{0}^{-1/2})\mathrm{vec}({\bf I}_{m})={\bf T}^{1/2}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})$. It follows from [51, rel. (14)] that
    ${\bm{\Pi}}^{\bot}_{\bm{\Delta}}={\bm{\Pi}}^{\bot}_{\bf V}-\frac{{\bm{\Pi}}^{\bot}_{\bf V}{\bf u}_{n}{\bf u}_{n}^{T}{\bm{\Pi}}^{\bot}_{\bf V}}{{\bf u}_{n}^{T}{\bm{\Pi}}^{\bot}_{\bf V}{\bf u}_{n}}.$ (212)
    Reporting expression (212) of ${\bm{\Pi}}^{\bot}_{\bm{\Delta}}$ in (210), condition (121) of Proposition VI.1 is satisfied iif
    $\mathrm{vec}({\bf I}_{m})^{T}{\bf T}^{-1/2}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k}-\frac{(\mathrm{vec}({\bf I}_{m})^{T}{\bf T}^{-1/2}{\bm{\Pi}}^{\bot}_{\bf V}{\bf u}_{n})({\bf u}_{n}^{T}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k})}{{\bf u}_{n}^{T}{\bm{\Pi}}^{\bot}_{\bf V}{\bf u}_{n}}=0,\ \ k=1,..,q.$ (213)
    Consequently to conclude the proof, it is sufficient to prove the two equalities:
    $\mathrm{vec}({\bf I}_{m})^{T}{\bf T}^{-1/2}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k}=0\ \ \mbox{and}\ \ {\bf u}_{n}^{H}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k}=0,\ \ k=1,..,q.$ (214)
    From the definition of ${\bf G}$ (208) and the derivative $\bm{\Sigma}_{k}^{0}={\bf A}_{k}^{0}\bm{\Xi}_{0}{\bf A}_{0}^{T}+{\bf A}_{0}\bm{\Xi}_{0}{({\bf A}_{k}^{0}})^{T}$, we straightforwardly deduce that
    ${\bf g}_{k}={\bf T}^{1/2}{\rm vec}({\bf Z}_{k}+{\bf Z}_{k}^{T})$ (215)
    with
    ${\bf Z}_{k}\triangleq\bm{\Sigma}_{0}^{-1/2}{\bf A}_{0}\bm{\Xi}_{0}{({\bf A}_{k}^{0}})^{T}\bm{\Sigma}_{0}^{-1/2}.$ (216)
    Likewise from the definition of ${\bf V}$ (211) and ${\bf J}_{{\rm vecs}(\bm{\Xi})}[{\rm vec}({\bf\Sigma}_{0})]=({\bf A}_{0}\otimes{\bf A}_{0}){\bf D}_{p}$, we straightforwardly get:
    ${\bf V}={\bf T}^{1/2}{\bf W}_{0}{\bf D}_{p}$ (217)
    with
    ${\bf W_{0}}\triangleq{\bf\Sigma}_{0}^{-1/2}{\bf A}_{0}\otimes{\bf\Sigma}_{0}^{-1/2}{\bf A}_{0}.$ (218)
    Consequently, ${\bm{\Pi}}^{\bot}_{\bf V}$ is written in the form:
    ${\bm{\Pi}}^{\bot}_{\bf V}={\bf I}_{m^{2}}-{\bf T}^{1/2}{\bf W}_{0}{\bf D}_{p}[{\bf D}_{p}^{T}({\bf W}_{0}^{T}{\bf T}{\bf W}_{0}){\bf D}_{p}]^{-1}{\bf D}_{p}^{T}{\bf W}_{0}^{T}{\bf T}^{1/2},$ (219)
    where ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}=\frac{\alpha(g_{0})}{2}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}\otimes{\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})+\frac{\alpha(g_{0})-1}{4}{\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0}){\rm vec}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})^{T}$. By noticing that ${\bf W}_{0}^{T}{\bf T}{\bf W}_{0}$ is structured in the form ${\bf A}+{\bf a}{\bf a}^{T}$ and applying the inversion matrix lemma with the trick in (186), we get:
    ${\bm{\Pi}}^{\bot}_{\bf V}={\bf I}_{m^{2}}-{\bf T}^{1/2}{\bf B}_{0}{\bf T}^{1/2},$ (220)
    with
    ${\bf B}_{0}\triangleq\left({\bf I}_{m^{2}}+{\bf K}_{m}\right)\left[\frac{1}{\alpha(g_{0})}({\bf H}_{1,0}\otimes{\bf H}_{1,0})-\frac{\alpha(g_{0})-1}{\alpha(g_{0})(2\alpha(g_{0})+(\alpha(g_{0})-1)p}{\rm vec}({\bf H}_{1,0}){\rm vec}({\bf H}_{1,0})^{T}\right]$ (221)
    with ${\bf H}_{1,0}\triangleq{\bf\Sigma}_{0}^{-1/2}{\bf A}_{0}({\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1}{\bf A}_{0})^{-1}{\bf A}_{0}^{T}{\bf\Sigma}_{0}^{-1/2}$. Finally, both left hand sides of expressions (214) follow
    $\mathrm{vec}({\bf I}_{m})^{T}{\bf T}^{-1/2}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k}$ $=$ $\mathrm{vec}({\bf I}_{m})^{T}\mathrm{vec}({\bf Z}_{k}+{\bf Z}_{k}^{T})-\mathrm{vec}({\bf I}_{m})^{T}{\bf B}_{0}{\bf T}\mathrm{vec}({\bf Z}_{k}+{\bf Z}_{k}^{T})$ (222)
    ${\bf u}_{n}^{H}{\bm{\Pi}}^{\bot}_{\bf V}{\bf g}_{k}$ $=$ $\mathrm{vec}({\bm{\Sigma}}_{0}^{-1})^{T}{\bf T}\mathrm{vec}({\bf Z}_{k}+{\bf Z}_{k}^{T})-\mathrm{vec}({\bm{\Sigma}}_{0}^{-1})^{T}{\bf T}{\bf B}_{0}{\bf T}\mathrm{vec}({\bf Z}_{k}+{\bf Z}_{k}^{T}).$ (223)
    Reporting the expressions of ${\bf T}$ (206), ${\bf Z}_{k}$ (216) and ${\bf B}_{0}$ (221) in the right hand side of (222) and (223), and using ${\bf H}_{1,0}^{2}={\bf H}_{1,0}$ and ${\rm tr}({\bf H}_{1,0})=p$, terms (222) and (223) are proven after cumbersome calculations to be equal to zero.